### v1

# The Galactic population of canonical pulsars II

Abstract
Context. Pulsars are highly magnetized rotating neutron stars, emitting in a broad electromagnetic energy range. These objects have been discovered more than 50 years ago and are astrophysical laboratories for studying physics at extreme conditions. Reproducing the observed population of pulsars refines our understanding of their formation, evolution as well as radiation processes and geometry.
Aims. In this paper, we improve our previous pulsar population synthesis by studying the impact of the Galactic gravitational potential and of the death line putting emphasis on the $\gamma$-ray pulsar population. In order to elucidate the necessity of a death line, refined initial distributions for spin periods and radial positions at birth were implemented, elevating the sophistication of our previous simulations to the most recent state-of-the-art.
Methods. The individual motion of pulsars in the Galactic potential is evolved by a fourth order symplectic integration scheme. Our pulsar population synthesis takes into account the secular evolution of a force-free magnetosphere and the magnetic field decay. Each pulsar is evolved from its birth up to the present time. The radio and $\gamma$-ray emissions are modeled respectively by the polar cap geometry and the striped wind model.
Results. When simulating a cohort of ten million pulsars akin to previous endeavors, the integration of a death line proves essential for a better agreement with observational trends. Yet, discrepancies emerges when modeling a reduced ensemble of only one million pulsars, endowed with a higher birth rate. This divergence yields a remarkably refined $P-\dot{P}$ diagram, conspicuously with or without any delineating death line. Intriguingly, this observation suggests an overestimation of the ages of detected pulsars in existent data sets, thereby challenging the necessity of death lines in pulsar population synthesis. Validation through Kolmogorov-Smirnov testing underlines the statistical congruence between the spin period and spin period derivative distributions of both observations and simulations. Moreover simulations with a higher $\gamma$-ray instrumental sensitivity show that the GeV excess in the Galactic centre could be of pulsar origin.
Key Words.:
pulsars: general - Gravitation - radio continuum: stars - Gamma rays: stars - methods: statistical
1 Introduction
[…]
Pulsars, discovered by Jocelyn Bell Burnell in 1967 (Hewish et al., 1968) are rotating neutron stars born in a core-collapsed supernova. Pulsars are highly magnetized and surrounded by a plasma-filled magnetosphere emitting regular pulses of radiation at their spin frequency. Due to the magneto-dipole losses, they lose energy as their spin period increases. Thanks to radio surveys, performed for instance by the Parkes and Arecibo radio-telescopes, nearly 3000 pulsars have been discovered (Manchester et al., 2005) (see Table 1, with the data on the canonical pulsars). Although pulsars were first observed in radio, it was found later that they are also bright in X-rays, optical and gamma-rays. The Large Area Telescope (LAT) on board of the Fermi satellite, has discovered dozens of radio-quiet gamma-ray pulsars as well as millisecond pulsars (MSPs). Since its launch, the LAT has detected about 300 pulsars (Abdo et al., 2013; Smith et al., 2019, 2023). The LAT has unveiled a new perspective, expanding our understanding about pulsars by presenting a growing collection of neutron stars that are detectable through their high energy emission. Moreover, the distinct emission process in high-energy and its unique beaming properties offer an alternative viewpoint on the global pulsar population.
Table 1: Number of known pulsars with spindown luminosity $\dot{E}$ above and below $10^{28}$ W, and above $10^{31}$ W. The quantities $N_{r}$, $N_{g}$, and $N_{rg}$ are the number of radio-only, gamma only, and radio-loud gamma-ray pulsars, respectively. It should be noted that we excluded the binary pulsars as well as pulsars with $P<20$ ms. The data have been taken from the ATNF catalogue and from https://confluence.slac.stanford.edu/display/ GLAMCOG/Public+List+of+LAT-Detected+Gamma-Ray+Pulsars
log($\dot{E}$) (W) $N_{tot}$ $N_{r}$ $N_{g}$ $N_{rg}$
¿31 2 0 0 2
¿28 197 101 35 61
Total 2665 2553 63 84
With the advent of future surveys such as the Square Kilometre Array (SKA  [^1] https://www.skao.int/en/science-users/118/ska-telescope-specifications) or the Cherenkov Telescope Array (CTA  [^2] https://www.cta-observatory.org/about/how-ctao-works/), much more pulsars are expected to be detected. Those instruments are scheduled to start collecting data by the end of the decade. SKA is going to be 50 times more sensitive and is predicted to survey the sky 10 000 times faster than any existing imaging radio telescope (McMullin et al., 2020). CTA is aimed at detecting gamma rays in the energy range from a few tens of GeV up to hundreds of TeV (Hofmann & Zanin, 2023) while Fermi/LAT operates in the energy range from 20 MeV to several hundred GeV  [^3] https://fermi.gsfc.nasa.gov/science/instruments/table1-1.html.
Pulsar population synthesis (PPS) is a powerful tool aiming at predicting the discovery rate of new pulsars as well as better understanding their emission processes. PPS studies also constrains the overall properties of pulsars. Pulsars are generated from their birth and evolved up to the present. Once the set of pulsars has been generated we check whether they fulfill some prescribed detection criteria. The detectability of an individual pulsar in a given energy band (radio or $\gamma$) depends on whether the associated emission beam intersects our line of sight and on the sensitivity of the corresponding instrument.
Most of PPS studies assume that neutron stars are rotating in vacuum (Faucher-Giguère & Kaspi, 2006; Popov et al., 2010; Johnston et al., 2020), or take only radio or gamma-ray emission into account (Watters & Romani, 2011; Gullon et al., 2014), or assume a constant magnetic field (Gonthier et al., 2002; Faucher-Giguère & Kaspi, 2006; Johnston & Karastergiou, 2017). For the first time in a PPS study, Dirson et al. (2022) have modeled the population of both the gamma-ray and radio pulsars by taking into account the force-free magnetosphere model in conjunction with the magnetic field decay. The neutron star radiation is produced by relativistic charged particles flowing within their magnetosphere. Therefore it is necessary to take the plasma back reaction into account for the pulsar period and dipole magnetic inclination angle evolution.Whereas the gamma-ray emission site from pulsars is still under debate, different locations have been suggested. These are for instance the polar cap region (Sturrock, 1971; Ruderman & Sutherland, 1975; Daugherty & Harding, 1982, 1996) at low altitudes, the slot gap along the last open magnetic field lines (Arons, 1983; Muslimov & Harding, 2004; Harding et al., 2008; Harding & Muslimov, 2011), or the outer gaps at high altitudes in the outer magnetosphere (Cheng et al., 1986; Hirotani, 2008; Takata et al., 2011). The striped wind model, which has for emission site the pulsar wind outside the light cylinder, (Pétri, 2009, 2011) is used for the first time by Dirson et al. (2022) in a PPS study. The radio emission is described as usual by the polar cap model.
In this work, and the work of Dirson et al. (2022) only the canonical pulsar population in our Galaxy is reproduced. The canonical pulsar population is defined by isolated pulsars which are not magnetars, and with a spin period greater than 20 ms.
While the study yielded satisfactory results when simulating a population closely resembling the observed pulsar population in the $P-\dot{P}$ diagram, there was room for improvement regarding the pulsar motion. Specifically, the study is limited by the assumption that pulsars move in straight lines at a constant speed from their birth to the present time. A noticeable discrepancy arises when comparing the observed position of pulsars to the simulation expectations, revealing a significant disparity in their spatial distributions.
Since the results obtained by Dirson et al. (2022) did not require the modeling of a death line, it was not included in their PPS. The death line separates in the $P-\dot{P}$ diagram the neutron stars that generate particles through pair cascading from those which cannot radiate anymore and are hence called dead pulsars. This contrasts with the PPS performed by Graber et al. (2023) who did not require a death line. Therefore in this paper we examine the necessity of a death line to explain the observed canonical pulsars population.
This paper improves the PPS of Dirson et al. (2022), questions the necessity of a death line in PPS work and discusses in depth the $\gamma$-ray pulsar population. The paper is organised as follows. In Sect. 2 we detail our model, recalling the processes for generation and evolution of the pulsars in the PPS, a description of the Galactic potential, the death line and the multi-wavelength detection criteria. Then the results are shown in Sect. 3. The paper ends with a discussion in Sect. 4 and a summary in Sect. 5.
2 Description of the PPS model
Our PPS study is based on a self-consistent state-of-the-art evolution of the pulsar geometry and proper motion within the Milky Way. In this section we detail the model by exposing first the generation of individual pulsars, their initial position, magnetic field and spin period, then how these quantities should evolve in time, giving explicit expressions for the Galactic potential used in our simulations, before introducing the death line and finally the radio and $\gamma$-ray detection criteria.
2.1 Generation of pulsars
We generate a population of pulsars whose ages are generated in ascending order, every X years for example if the birth rate is 1/X yr-1 between 0 yr and the age of the Milky Way at maximum. The age of the Milky Way is almost equal to the age of the Universe, which is of about $t_{H}\approx 13.8\times 10^{9}$ yr, therefore, a total number of $t_{H}/X$ pulsars should be simulated. However as we will see, we chose to emulate $10^{7}$ pulsars in a first run and $10^{6}$ in another run. This choice was made for two reasons : firstly, the computation time is significantly decreased while pulsars older than $10^{8}$ yr are mostly not detected, secondly, the simulation that emulates $10^{7}$ pulsars allows to emulate older pulsars and require a death line while the other works well without. This approach was also chosen because it allows to check whether our birth rate is relevant by comparing our results from the simulation to the observations. It is different from other works like Johnston & Karastergiou (2017) and Gullon et al. (2014), where they stop the simulation when the number of pulsars detected in the simulation are the same as the number of observed pulsars. The value chosen for the birth rate, which is 1/41 yr-1, is taken as best educated guess from the interval [1/150,1/33] yr-1 because these studies (Faucher-Giguère & Kaspi, 2006; Gullon et al., 2014; Johnston & Karastergiou, 2017) evaluated the birth rate of pulsars in this interval.
To describe the position of the pulsars in the Galaxy, we use the right-handed Galacto-centric coordinate system $(x,y,z)$ with the Galactic centre at its origin, $y$ increasing in the disc plane towards the location of the Sun, and $z$ increasing towards the direction of the north Galactic pole. The initial spatial distribution of the pulsars is given by Paczynski (1990). He found two distributions, one describing the radial spread and one the spread in altitude, however only his distribution in altitude is used in our work because the radial distribution in the galactic plane is better depict in Ahlers et al. (2016) by the Milky Way’s pulsar surface density defined as
$\rho(R)$ $=A\left(\frac{R+R_{1}}{R_{\odot}+R_{1}}\right)^{a}\exp\left(-b\left(\frac{R-R_{\odot}}{R_{\odot}+R_{1}}\right)\right)$ (1a)
$\rho_{z}(z)$ $=\frac{e^{-|z|/h_{c}}}{h_{c}}$ (1b)
where $R$ is the axial distance from the $z$-axis, and $z$ is the distance from the Galactic disc. The numerical values for the constants are, $A=37.6$ kpc-2, $a=1.93$, $b=5.06$, $R_{1}=0.55$ kpc, $h_{c}=180$ pc and $R_{\odot}=8.5$ kpc and in agreement with the distribution of young massive stars in our galaxy (Li et al., 2019). The purpose of choosing equation (1a) to describe the radial spread is to put neutron stars birth positions within the spiral arms of the Galaxy. More precisely the Galactic spiral structure contains four arms with a logarithmic shape function allowing to obtain the azimuthal coordinate $\phi$ as a function of the distance from the Galactic center
$\phi(R)=k\ln\left(\frac{R}{r_{0}}\right)+\phi_{0}$ (2)
The values of the models describing each arm are given in Table 2. These values were used by Ronchi et al. (2021) and taken from Yao et al. (2017) to match the shape of the arms of the Galaxy. The local arm is not modeled, because of its very low density, much smaller than the four other arms. Each star has an equal probability to be in one of the four arms, its angular coordinate $\phi$ for a given $R$ being deduced from (1a). Furthermore, the Galaxy is not static, and its arms are moving with an approximated period $T=250$ Myr (Skowron et al., 2019). Since we also know that the Galaxy is rotating in the clockwise direction, by knowing the age of a pulsar we can know its angular position at birth. Following the same procedures as Ronchi et al. (2021), in order to avoid artificial features near the Galactic center, noise is added to both coordinates R and $\phi$. For instance $\phi_{corr}=\phi_{rand}\,\exp(-0.35\,R)$ where $\phi_{rand}$ is randomly drawn from a uniform distribution between 00 and $2\,\pi$. $r_{corr}$ will be taken from a normal distribution with a mean of 00 and a standard deviation $\sigma_{corr}=0.07\,R$ and we add this two corrections to $\phi$ and $R$ respectively. Therefore $R_{birth}=R+r_{corr}$ and $\phi_{birth}=\phi+\phi_{corr}+\frac{2\,\pi\,t_{age}}{T}$, where $t_{age}$ is the age of the pulsar. As soon as $R_{birth}$ and $\phi_{birth}$ are known we can convert these coordinates in $x$ and $y$ Galacto-centric coordinates for each pulsar.
Table 2: Parameters of the Milky Way Spiral Arm structure.
Arm Number Name $k$ $r_{0}$ $\phi_{0}$
(rad) (kpc) (rad)
1 Norma 4.95 3.35 0.77
2 Carina-Sagittarius 5.46 3.56 3.82
3 Perseus 5.77 3.71 2.09
4 Crux-Scutum 5.37 3.67 5.76
Next, we need to describe the individual properties of each pulsar, namely, its inclination angle $\alpha$, which is the angle between its rotation axis and its magnetic axis, its initial spin period and initial magnetic field at birth. The inclination angle $\alpha$ is assumed to follow an isotropic distribution generated from a uniform distribution $U\in[0,1]$ and given by $\alpha=\arccos(2\,U-1)$. In our work, the initial spin period and magnetic field follow both a log-normal distribution, as suggested from the results of a study of 56 young neutron stars by Igoshev et al. (2022) while other works simply use a Gaussian distribution (Faucher-Giguère & Kaspi, 2006; Gullon et al., 2014; Johnston et al., 2020). Explicitly, the probabilities to born with an initial magnetic field $B_{0}$ and period $P_{0}$ are given by
$p(\log(B_{0}))=\frac{1}{\sigma_{b}\sqrt{2\pi}}e^{-(\log B_{0}-\log\bar{B})^{2}/(2\sigma_{b}^{2})}$ (3)
$p(\log(P_{0}))=\frac{1}{\sigma_{p}\sqrt{2\pi}}e^{-(\log P_{0}-\log\bar{P})^{2}/(2\sigma_{p}^{2})}.$ (4)
The distribution for the magnetic field is identical to the one used in other pulsar population synthesis like for instance Faucher-Giguère & Kaspi (2006), Gullon et al. (2014), Johnston et al. (2020), Yadigaroglu & Romani (1995) and Watters & Romani (2011).
Compared to the study of Dirson et al. (2022) we introduce a novelty, namely the alignment between the kick velocities at birth and the rotation axis of the pulsar, as suggested by Rankin (2007). Moreover according to Hobbs et al. (2005) a Maxwellian distribution best replicates the observations, with a standard deviation $\sigma_{\rm v}=265$ km/s. The mean velocity of the distribution is related to the standard deviation by $\bar{v}=\sigma_{\rm v}\sqrt{{8}/{\pi}}$. The velocity vector is then distributed along the unit vector of the rotation axis generated as detailed in the subsection 2.5.
2.2 Pulsar evolution
The pulsar initial period, inclination angle and magnetic field must be evolved in time self-consistently due to electromagnetic radiation and internal dissipation within the neutron star. To this end we employ the pulsar force-free magnetosphere model taking into account the plasma current and charge within the magnetosphere. The spin down luminosity $\dot{E}$ is
$\dot{E}=\frac{dE_{\rm rot}}{dt}=-I\Omega\dot{\Omega}=L_{\rm ffe}$ (5)
$L_{\rm ffe}=\frac{4\pi R^{6}B^{2}\Omega^{4}(1+\sin^{2}\alpha)}{\mu_{0}c^{3}}$ (6)
where $I$ is the neutron star moment of inertia its value being approximately $I\approx 10^{38}$ kg.m2, $\Omega=2\pi/P$ is the rotation frequency of the pulsar ($P$ is the spin period) and $\dot{\Omega}$ is its time derivative, $R=12$ km is the typical radius of a neutron star as found by recent NICER observations (Riley et al., 2019; Bogdanov et al., 2019), $B$ is the magnetic field at the magnetic equator, $\alpha$ is the inclination angle, $\mu_{0}$ is the vacuum permeability constant, $\mu_{0}=4\pi\times 10^{-7}$ H/m and $c$ is the speed of light. The term $L_{\rm ffe}$ for the spin down luminosity was given by Spitkovsky (2006) and Pétri (2012). Equation (5) clearly shows the correlation between the obliquity, the magnetic field and the rotation frequency. Combining equation (5) and (6) leads to
$\dot{\Omega}=-K_{\rm ffe}\ \Omega^{n}$ (7)
$K_{\rm ffe}=\frac{4\pi R^{6}B^{2}(1+\sin^{2}\alpha)}{\mu_{0}c^{3}I}$ (8)
which is written in a more concise form where $n$ is the braking index, its value being $n=3$ for magnetic dipole radiation.
The integral of motion between $\Omega$ and $\alpha$ is also very useful and reads
$\Omega\frac{\cos^{2}\alpha}{\sin\alpha}=\Omega_{0}\frac{\cos^{2}\alpha_{0}}{\sin\alpha_{0}}$ (9)
where quantities with index 0 indicate their initial value. The quantities without the index are their current time value.
As shown by Gullon et al. (2014) a decaying magnetic field better accounts for comparison between pulsar population synthesis and observations. Therefore, we prescribe a magnetic field decay according to a power law
$B(t)=B_{0}(1+t/\tau_{d})^{-1/\alpha_{d}}$ (10)
where $\alpha_{d}$ is a constant parameter controlling the speed of the magnetic field decay and $\tau_{d}=k_{\tau_{d}}\tau_{v}$ is the magnetic field decay time scale. The decay rate $\tau_{v}$ was chosen randomly between 3 values : $5\times 10^{4}$ yr or $7\times 10^{4}$ yr or $5\times 10^{5}$ yr, with a probability of 0.23, 0.46 and 0.31 respectively (these probabilities were chosen as best educated guesses after several run with different values) for each pulsar, corresponding to the $\tau_{v}$ value for canonical pulsars in Viganò et al. (2013). The values of $\alpha_{d}$ and $k_{\tau_{d}}$ can be found in Table 5 of Sect. 3.
In line with the spin down luminosity, the inclination angle satisfies another evolution equation that after integration was found by Phillipov et al. (2014) for a spherically symmetric neutron star with a constant magnetic field. With our decaying prescription the inclination angle $\alpha$ is found by solving for the root of
$\begin{split}\ln(\sin\alpha_{0})+\frac{1}{2\sin^{2}\alpha_{0}}+K\Omega_{0}^{2}\frac{\cos^{4}\alpha_{0}}{\sin^{2}\alpha_{0}}\frac{\alpha_{d}\tau_{d}B_{0}^{2}}{\alpha_{d}-2}\left[\left(1+\frac{t}{\tau_{d}}\right)^{1-2/\alpha_{d}}-1\right]\\ \phantom{{}={}}=\ln(\sin\alpha)+\frac{1}{2\sin^{2}\alpha}\end{split}$ (11)
where $t$ is the time representing the age of the pulsar and $K={R^{6}}/{I\,c^{3}}$. The typical decay timescale for a mean magnetic field of $2.5\times 10^{8}$ T is $4.6\times 10^{5}$ yr in Viganò (2013), however in this study the typical decay timescale associated to this magnetic field is $3.5\times 10^{5}$ yr which is consistent with the estimate of Viganò (2013) for the same field strength.
Finally, we need to follow the particle motion within the Galactic potential. Each generated pulsar evolves in the gravitational potential $\Phi$ subject to an acceleration ${\ddot{x}}$ according to
${\ddot{x}}=-\@vec{\nabla}\Phi\ .$ (12)
Equation (12) is integrated numerically thanks to a Position Extended Forest Ruth-Like (PEFRL) algorithm (Peter Young, 2014), a fourth order integration scheme, presented in Appendix A with a convergence and accuracy study. The galactic potential model used in discussed in the next section.
2.3 Description of the Galactic potential
The Galaxy is divided in four distinct regions with different mass distributions and associated gravitational potentials. The four potentials are: the bulge $\Phi_{b}$, the disk $\Phi_{d}$, the dark matter halo $\Phi_{h}$ and the nucleus $\Phi_{n}$. The total potential of the Milky Way $\Phi_{tot}$ is the sum of these potentials
$\Phi_{tot}=\Phi_{b}+\Phi_{d}+\Phi_{h}+\Phi_{n}.$ (13)
The expressions for these potentials are taken from Bajkova & Bobylev (2021) with parameters given in Table 3. The nucleus mass was found in Bovy (2015).
Table 3: Constants values used for the different potentials, with the solar mass $M_{\odot}=1.99\times 10^{30}$ kg.
Parameters Values
$M_{h}$ $2.9\times 10^{11}\pm 7.6\times 10^{10}M_{\odot}$
$a_{h}$ 7.7 $\pm$ 2.1 kpc
$a_{d}$ 4.4 $\pm$ 0.73 kpc
$b_{d}$ 0.308 $\pm$ 0.005 kpc
$M_{d}$ $6.50\times 10^{10}\pm 1.9\times 10^{9}M_{\odot}$
$a_{b}$ 0.0 kpc
$b_{b}$ 0.267 $\pm$ 0.009 kpc
$M_{b}$ $1.02\times 10^{10}\pm 6.3\times 10^{8}M_{\odot}$
$M_{n}$ $4.10^{6}\pm 0.42\times 10^{6}M_{\odot}$
The potential for the bulge $\Phi_{b}$ and for the disk $\Phi_{d}$ were both chosen to have the form proposed by Miyamoto & Nagai (1975) which are typically used for models of the gravitational potential of the Milky Way. They read
$\Phi_{i}(R,z)=-\frac{GM_{i}}{\left[R^{2}+\left(a_{i}+\sqrt{z^{2}+b_{i}^{2}}\right)^{2}\right]^{1/2}}$ (14)
where $R^{2}=x^{2}+y^{2}$, and $i=b$ is for the bulge and $i=d$ is for the disk. In those formula $R$ depends on the coordinates $x$ and $y$, $a_{i}$ and $b_{i}$ are the scale parameters of the components in kpc, $M_{i}$ is the mass (for the disk or the bulge) and $G$ is the gravitational constant. The constant values can be found in Table 3.
Solving the pulsar equation of motion requires to compute the gradient of the potentials, the derivatives of the disk and the bulge potentials are
$\frac{\partial\Phi_{i}}{\partial x_{j}}=\frac{GM_{i}x_{j}}{\left[R^{2}+\left(a_{i}+\sqrt{z^{2}+b_{i}^{2}}\right)^{2}\right]^{3/2}}$ (15)
$\frac{\partial\Phi_{i}}{\partial z}=\frac{GM_{i}z\left(a_{i}+\left(z^{2}+b_{i}^{2}\right)^{\frac{1}{2}}\right)}{\left[R^{2}+\left(a_{i}+\sqrt{z^{2}+b_{i}^{2}}\right)^{2}\right]^{3/2}\ \left(z^{2}+b_{i}^{2}\right)^{1/2}}$ (16)
$x_{j}$ being either $x$ or $y$. Equation (15) is the derivative for the potential of Miyamoto Nagai, where $i=d$ or $i=b$ depending on disk or bulge derivative. Equation (16) is the derivative of the Miyamoto Nagai potential with respect to the $z$ coordinate.
The potential for the dark matter halo $\Phi_{h}$, according to the frequently used potential of Navarro et al. (1997), it is expressed as
$\Phi_{h}(r)=-\frac{GM_{h}}{r}\ \ln\left(1+\frac{r}{a_{h}}\right)$ (17)
where $M_{h}$ is the mass of the halo, $r=x^{2}+y^{2}+z^{2}$, and $a_{h}$ is the length scale whose values are found in Table 3. The derivative of equation (17) with respect to any coordinate $x_{j}$ $(x,y,z)$ is
$\frac{\partial\Phi_{h}}{\partial x_{j}}=\frac{GM_{h}x_{j}}{r^{2}}\ \left[\frac{1}{r}\ \ln\left(1+\frac{r}{a_{h}}\right)-\frac{1}{r+a_{h}}\right].$ (18)
Finally for the last part of the Galaxy, the nucleus of the Milky Way is simply represented by a Keplerian potential $\Phi_{n}$ such as
$\Phi_{n}(r)=-\frac{GM_{n}}{r}$ (19)
where $M_{n}$ is the mass of the nucleus. The derivative of equation (19) with respect to any coordinate $x_{j}\ (x,y,z)$ is given by
$\frac{\partial\Phi_{n}}{\partial x_{j}}=\frac{GM_{n}\,x_{j}}{r^{3}}.$ (20)
2.4 The death line
Several equations for the death line have been proposed in the literature (Chen & Ruderman, 1993; Zhang et al., 2000; Gil & Mitra, 2001; Mitra et al., 2019). The most suitable one in our study is that of Mitra et al. (2019) where they found an expression for the death line
$\dot{P}_{\rm line}=\frac{3.16\times 10^{-19}\ T_{6}^{4}\ P^{2}}{\eta^{2}b\cos^{2}\alpha_{l}}\ .$ (21)
Here $T_{6}=T/10^{6}$ K, where $T$ corresponds to the surface temperature of the polar cap. The parameter $\eta=1-\rho_{i}/\rho_{GJ}$ is the electric potential screening factor due to the ion flow, where $\rho_{i}$ correspond to the ion charge density and $\rho_{GJ}$ to the Goldreich-Julian charge density above the polar cap. The quantity $b$ is the ratio of the actual surface magnetic field to the dipolar surface magnetic field and $\alpha_{l}$ is the angle between the local magnetic field and the rotation axis. The parameters we used for this death line are : $\alpha_{l}=45\degr$, $b=40$, $\eta=0.15$, $T_{6}=2$. Spread in the parameter values of the model causes significant variations in the death line, and thus a death valley rather than a single death line describes the condition for pulsar extinction in the $P-\dot{P}$ diagram.
To decide whether a pulsar is still active in radio, its critical spin down rate $\dot{P}_{\rm line}$ is computed knowing its spin period $P$. If the pulsar lies in the death valley defined by the inequality $10^{-0.55}<\dot{P}_{\rm line}/\dot{P}<10^{1.15}$ then a value for $\alpha_{l}$ is taken from a uniform distribution between $0\degr$ and $65\degr$. Moreover a value for $T_{6}$ is also taken from a uniform distribution between 1.9 and 2.8. The parameter $b$ is taken from a uniform distribution between 30 and 60. The boundary parameter values for the death valley were chosen to traverse the $P-\dot{P}$ diagram from the region in the bottom right quadrant where the density starts to decrease, to the detectable pulsar with the lowest $\dot{P}$ and highest $P$ in the same quadrant, as shown in Fig. 1. The edges of the death valley given by the values $10^{-0.55}$ and $10^{1.15}$ were found by taking the extreme values for the parameters that are drawn from the different uniform distributions. Then we compare its actual spin down rate $\dot{P}$, computed from the evolution model, to the $\dot{P}^{\rm new}_{\rm line}$ computed with equation (21) with the parameters which were drawn from uniform distributions. For pulsars that are not in the death valley, their $\dot{P}$ is only compared with the spin down rate $\dot{P}_{\rm line}$ obtained with the default parameters. If the pulsar has its $\dot{P}$, which was computed from the evolution model, below the one computed with the death line ($\dot{P}_{\rm line}$ or $\dot{P}^{\rm new}_{\rm line}$), then the pulsar is considered as dead, otherwise it is considered as still emitting photons. These choices were made in order to have more like a death valley near the death line than a strict condition of death, because the parameters $T_{6}$, $\alpha_{l}$ and $b$ may be different from one pulsar to another.
Figure 1: $P-\dot{P}$ diagram of the canonical pulsars along with the death line, green solid line, and death valley, shaded green area.
2.5 Detection
For each generated pulsar, we check whether it fulfills the criteria for detection depending on three factors. First, the beaming fraction indicates the fraction of the sky covered by the radiation beam and depends on the considered wavelength, here radio or $\gamma$-ray. The beaming fraction varies also with the pulsar spin rate, geometry and location of the emission regions. Before explaining how the beaming fraction in radio and in $\gamma$-ray are computed, it is important to define several angles.
The angle between the line of sight and the rotation axis is denoted by $\xi=(\widehat{\@vec{n_{\Omega}},\@vec{n}})$ where ${n}_{\Omega}={\@vec{\Omega}}/{\|\@vec{\Omega}\|}$ is a unit vector along the rotation axis and ${n}$ the unit vector along the line of sight. The inclination angle $\alpha=(\widehat{\@vec{n_{\Omega}},\@vec{\mu}})$ is the angle between the rotation axis and the magnetic moment, ${\mu}$ being the unit vector along the magnetic moment. Traditionally the impact angle is also introduced as the angle between the magnetic moment and the line of sight $\beta=(\widehat{\@vec{\mu},\@vec{n}})$. Moreover, it is related to the previous angles by $\alpha+\beta=\xi$.
We choose a isotropic distribution for the Earth viewing angle $\xi$ as well as for the orientation of the unitary rotation vector. The Cartesian coordinates of the unit rotation vector ${n_{\Omega}}$ are $(\sin\theta_{n_{\Omega}}\cos\phi_{n_{\Omega}},\sin\theta_{n_{\Omega}}\sin\phi_{n_{\Omega}},\cos\theta_{n_{\Omega}})$. We set the Sun position at $(x_{\odot},y_{\odot},z_{\odot})=$ (0 kpc, 8.5 kpc, 15 pc) (Siegert, 2019). The coordinates for ${n}$ are
${n}=\left(\frac{x-x_{\odot}}{d},\frac{y-y_{\odot}}{d},\frac{z-z_{\odot}}{d}\right).$ (22)
To compute the pulsar distance from Earth, we use the formula for the distance which is
$d=\sqrt{(x-x_{\odot})^{2}+(y-y_{\odot})^{2}+(z-z_{\odot})^{2}}.$ (23)
2.5.1 Radio detection model
The beaming fraction in radio depends on the half opening angle of the radio emission cone $\rho$ computed according to
$\rho=3\sqrt{\frac{\pi\,h_{\rm em}}{2Pc}}$ (24)
where $h_{\rm em}$ is the emission height, $P$ is the spin period of the pulsar and $c$ is the speed of light. The emission height is taken constant with an average value of $h_{\rm em}=3.10^{5}$ m, estimated from observations of a sample of pulsars by (Weltevrede & Johnston, 2008; Mitra, 2017; Johnston & Karastergiou, 2019; Johnston et al., 2023). The cone half-opening angle $\rho$ estimated in eq. (24) holds only for the last open field lines of a magnetic dipole and can only be applied for slow pulsars where the radio emission altitude is high enough for the multipolar components to decrease significantly and become negligible.
The pulsar is detected in radio if $\beta=|\xi-\alpha|\leq\rho$ or if $\beta=|\xi-(\pi-\alpha)|\leq\rho$ corresponding to the north and south hemisphere respectively. It must also satisfy the condition $\alpha\geq\rho$ and $\alpha\leq\pi-\rho$ in order to effectively see radio pulsation, because the line of sight must cross the emission cone to observe pulsations. Another useful quantity is the observed width of the radio profile $w_{r}$ which is computed as (see Lorimer & Kramer (2004))
$\cos(\rho)=\cos(\alpha)\cos(\xi)+\sin(\alpha)\sin(\xi)\cos(w_{r}/2).$ (25)
The second factor for detection is the luminosity, which is also different between radio and gamma-rays. First, concerning the radio flux density, the formula used is the same as in Johnston et al. (2020) for a pulsar at 1.4 GHz, in order to model the detection carried out by the Parkes radio telescope in the Southern Galactic plane (Kramer et al., 2003; Lorimer et al., 2006; Cameron et al., 2020) and the Arecibo telescope in the northern plane (Cordes et al., 2006)
$F_{r}=9\ \text{mJy}\left(\frac{d}{1\ \text{kpc}}\right)^{-2}\left(\frac{\dot{E}}{10^{29}\ \text{W}}\right)^{1/4}\times 10^{F_{\rm j}}$ (26)
where $d$ is the distance in kpc and $F_{\rm j}$ is the scatter term which is modeled as a Gaussian with a mean of 0.0 and a variance of $\sigma$ = 0.2. The detection threshold in radio is set by the signal to noise ratio defined by
$S/N=\frac{F_{r}}{S_{\rm survey}^{\rm min}}.$ (27)
The pulsar is detected if the signal to noise ratio $S/N$ is greater or equal to 10. $S_{\rm survey}^{\rm min}$ is the minimum flux which is related to the period of rotation $P$, the width profile of radio emission $w_{r}$ and the sensitivity of the survey which is the last factor for detection, and will be detailed later. We directly compute the radio flux, without computing the luminosity in radio which overlooks the fact that the luminosity received will depend on the geometry of the beam.
The third and last factor is the sensitivity, depending on the survey, therefore on the instrument used and on the pulse profile observed. A pulsar is detected in radio if the signal to noise ratio is greater than 10, but this ratio depends on the minimum flux, another function of the instrumental sensitivity. With the aim of computing the observed width pulse profile, we use the same formula as Cordes & McLaughlin (2003)
$\tilde{w}_{r}=\sqrt{\left(w_{r}\frac{P}{2\pi}\right)^{2}+\tau_{samp}^{2}+\tau_{DM}^{2}+\tau_{scat}^{2}}$ (28a)
$S_{\text{survey}}^{\text{min}}=S_{0}\sqrt{\frac{\tilde{w}_{r}}{P-\tilde{w}_{r}}}$ (28b)
Equation (28a) takes into account density fluctuations in the interstellar medium (ISM), the dispersion and scattering during the propagation of the radio pulse when interacting with the free electrons. Moreover, the instrumental effect is also taken into account with $\tau_{samp}$, which is the sampling time of the instrument. All this mechanisms lead to a broader pulse profile. In order to determine the influence of the ISM through $\tau_{DM}$, the formula of Bates et al. (2014) is used
$\tau_{DM}=\frac{e^{2}\Delta f_{ch}\,DM}{\pi\,m_{e}\,c\,f^{3}}$ (29)
$e$ is the electron charge, $m_{e}$ its mass, $\Delta f_{ch}$ the width of frequency of the instrument channel, $f$ the observing frequency and $DM$ the dispersion measure. Furthermore, in order to determine the influence of scattering by an inhomogeneous and turbulent ISM through $\tau_{scat}$, the empirical fit relationship from Krishnakumar et al. (2015) is used
$\tau_{scat}=3.6\times 10^{-9}DM^{2.2}(1+1.94\times 10^{-3}DM^{2})\ .$ (30)
Both $\tau_{DM}$ and $\tau_{scat}$ depend on the dispersion measure, which is found for each pulsar by running the code of Yao et al. (2017) converting the distance of a pulsar into the dispersion measure thanks to their state-of-the art model of the Galactic electron density distribution. The radio survey parameters used are the ones from the Parkes Multibeam Pulsar Survey (PMPS), see Table 4.
Table 4: Survey parameters of the Parks Multibeam Pulsar Survey (PMPS).
$\Delta f_{ch}$ $f$ $\tau_{samp}$
(kHz) (GHz) (µs)
3000 1.374 250
Concerning equation (28b), $S_{0}$ represents the survey parameters. Johnston et al. (2020) estimated that $S_{0}$ should approximately be equal to 0.05 mJy to have a signal to noise ratio greater than 10 for normal pulsars in this survey. We chose equation (28b) to compute the minimum flux, to decide whether a pulsar will be detected in radio or not, however it does not reproduce the whole complexity of this quantity which could be computed more precisely, (see equation (24) of Faucher-Giguère & Kaspi (2006) for instance, where $S_{0}$ is computed more precisely) even though it remains a good approximation. In addition we only use the parameters of PMPS to compute $\tau_{DM}$, we could have also used the parameters of The Pulsar Arecibo L-band Feed Array (PALFA) Survey, however the differences in the results in the end are not large if we use these parameters, consequently, it persists as a favorable approximation to exclusively utilize the parameters delineated by PMPS.
2.5.2 Gamma-ray detection model
The $\gamma$-ray emission model relies on the striped wind model, describing gamma-ray photons production emanating from the current sheet within the striped wind. In order to detect the gamma-ray, the line of sight of the observer must remain around the equator plane with an inclination angle constrained by $|\xi-\pi/2|\leq\alpha$. The gamma-ray luminosity is extracted from a study of Kalapotharakos et al. (2019), where they showed that the luminosity is described by a fundamental plane. The 3D model depends on the magnetic field $B$, the spin down luminosity $\dot{E}$ and the cut-off energy $\epsilon_{cut}$. However, in our PPS the cut-off energy is not computed therefore we use their two dimensional version
$L_{\gamma(2D)}=10^{26.15\pm 2.6}\ W\ \left(\frac{B}{10^{8}\ \text{T}}\right)^{0.11\pm 0.05}\ \left(\frac{\dot{E}}{10^{26}\ \text{W}}\right)^{0.51\pm 0.09}$ (31)
The spindown $\dot{E}$ is computed with equation (5) and the associated $\gamma$-ray flux detected on earth is computed with
$F_{\gamma}=\frac{L_{\gamma(2D)}}{4\,\pi\,f_{\Omega}\,d^{2}}$ (32)
where $f_{\Omega}$ is a factor depending on the emission model reflecting the anisotropy. For the striped wind model Pétri (2011) showed that this factor varies between 0.22 and 1.90. Nevertheless, an approximation is made : if $\alpha<-\xi+0.6109$ then $f_{\Omega}=1.9$ otherwise $f_{\Omega}=1$. The pulsar is detected in gamma depending on the instrumental sensitivity as described below.
The sensitivity in gamma-ray is based on the expectation of the Fermi/LAT instrument  [^4] https://fermi.gsfc.nasa.gov/ssc/data/analysis/documentation/Cicerone
/Cicerone_LAT_IRFs/LAT_sensitivity.html
[…]
3 Results
Let us now move to the PPS simulation results. First we show runs simulating 10 million pulsars with and without a death line, allowing to simulate old pulsars (with the oldest aged $4.1\times 10^{8}$ yr). Then we show runs simulating 1 million pulsars (with the oldest aged $4.1\times 10^{7}$ yr) with and without considering a death line. We explore a possible spin-velocity misalignment effect due to the galactic potential and the impact of the initial spin period distribution, normal vs log-normal. Finally, the gamma-ray simulations and detection are discussed in details.
3.1 Simulating 10 millions pulsars
3.1.1 Without a death line
The first simulations performed with the Galactic potential implemented discarded the death line. The parameters used for these runs are given in Table 5. The total number of pulsars is summarized in Table 6.
Table 5: Parameters used in the simulations.
$\tau_{\text{birth}}$ (1/yr) $P_{\text{mean}}$ (ms) $B_{\text{mean}}$ (T) $\sigma_{p}$ $\sigma_{b}$ $\alpha_{d}$ $k_{\tau_{d}}$
41 129 $2.75\times 10^{8}$ 0.45 0.5 1.5 5
Table 6: Number of pulsars detected without the implementation of the death line.
log($\dot{E}$) (W) $N_{tot}$ $N_{r}$ $N_{g}$ $N_{rg}$
¿31 11 0 8 3
¿28 161 9 75 77
Total 2653 2336 139 178
The quantities $N_{r}$, $N_{g}$, and $N_{rg}$ are the number of radio-only, gamma-only, and radio-loud gamma-ray pulsars, respectively, obtained from our simulation.
Compared to the results obtained by Dirson et al. (2022), at very high spin down luminosity, $\dot{E}>10^{31}$ W, the detection is almost identical but at $\dot{E}>10^{28}$ W there are much less pulsars detected in radio and radio/gamma in this work. There are probably two causes for this discrepancy in the number of detected pulsars at $\dot{E}>10^{28}$ W : firstly the difference in the spin period distribution at birth. Pulsars have a higher probability to start with a higher $P_{0}$ since the mean of the distribution is higher (in the previous study the mean was 60 ms while it is 129 ms now), therefore the slowing down of the pulsars can only decrease their rotation frequency $\Omega$, and equation (5) shows that $\dot{E}$ depends on $\Omega$. However, it does not affect the detection in gamma-rays because usually canonical gamma pulsars have a high $\dot{P}$ and a low $P$, being young, therefore their $\dot{E}$ is higher than for radio pulsars. Secondly, the ISM dispersion and scattering reduces this number of radio pulsars detected (the results without the ISM effect are not shown but without it there is much more radio pulsars detected). Nonetheless, in terms of total number of detected radio pulsars, the Galactic potential allows to have an increase compared to the number found in Dirson et al. (2022), a result expected because of the attractive nature of the potential of the Galaxy. In its previous version the PPS moved the pulsars in random spatial directions at constant speed, as a result no pulsar was bound to the Milky Way in closed orbits this is especially true for old pulsars which could have time to leave the Galaxy.
Figure 2: Distribution of the observed period taken from the ATNF catalogue, along with the simulations without the implementation of the death line. In green, the period of the simulated pulsars with age greater than $10^{7.5}$
Fig. 2 shows the distribution of the spin period of the detected pulsars in the simulation, in red and in observations, in blue. However the comparison is not satisfactory because there is an excess of pulsars detected in the simulation. Furthermore, this excess is also present in Fig. 3 where the age of the detected pulsars from the observations and simulation are plotted.
In order to verify if there is a correlation between old pulsars and the ones with a high spin period, we plotted in Fig. 2, in green the pulsars with an age greater than $10^{7.5}$ yr, corresponding to the oldest pulsars of the simulation and it corresponds to the excess found in Fig. 2. Due to the Galactic potential, more pulsars remain bound to the Galaxy, especially older pulsars which were escaping the Galaxy more easily in the work of Dirson et al. (2022) because of the absence of potential. Fig. 4 shows that many pulsars are in what is commonly called the graveyard of pulsars at the bottom right of the $P-\dot{P}$ diagram.
Figure 3: Distribution of the observed age taken from the ATNF catalogue, along with the simulations without the implementation of the death line.
Figure 4: $P-\dot{P}$ diagram of the simulated population, along with the observations, without the implementation of the death line.
Given this excess of long period pulsars without a death line, we decided to implement one, in order to stick closer to reality.
3.1.2 With the death line
Adding a death line decreases the number of pulsars detected in radio compared to the situation without a death line, see Table 7.
Table 7: Number of pulsars detected with the death line implementation.
log($\dot{E}$) (W) $N_{tot}$ $N_{r}$ $N_{g}$ $N_{rg}$
¿31 10 0 5 5
¿28 118 7 69 42
Total 2077 1807 116 154
Figure 5: Distribution of the observed period taken from the ATNF catalogue, along with the simulations with the death line implementation.
Figure 6: $P-\dot{P}$ diagram of the simulated population, along with the observations, with the death line implementation.
Figure 7: Density plot of the $P-\dot{P}$ diagram in comparison with observations
The death line improves the fit as seen in Fig. 5 for the distribution of spin period and compared with Fig. 2, even though there is a slight excess of pulsars which have spin period between $10^{-0.5}$ s and $1$ s. In Fig 6, these pulsars in excess lie at $\dot{P}$ between $10^{-18}$ and $10^{-17}$, where there is no data and which correspond to the oldest pulsars in the simulation. Therefore the oldest pulsars simulated here do not correspond to any real data.
In addition, there are too many pulsars close to the death line compared to the observations, there is a pile up effect created by the death line which is not in the observations.
The goodness-of-fit can be checked using a density plot of the $P-\dot{P}$ diagram as in Johnston & Karastergiou (2017) instead of putting points for individual pulsars. In order to construct such diagrams we binned the data evenly in a log scale in $P$ and $\dot{P}$ and count the number of pulsars in each bin. This leads to a 2D histogram to be compared with the observed density plot 2D histogram. Because the total number of simulated pulsars $N_{totsim}$ is different from the total number of observed pulsars $N_{totobs}$, we normalise the number of pulsars in each bin in order to compare proportions. We therefore introduce the quantity $R$ as follows
$R=\frac{N_{sim}/N_{totsim}-N_{obs}/N_{totobs}}{N_{sim}/N_{totsim}+N_{obs}/N_{totobs}}.$ (33)
With this definition the number $R$ remains in the interval $[-1,+1]$. $N_{sim}$ is the number of pulsar count in a bin of the simulation, $N_{obs}$ the number of pulsar detected in a bin of the observations. A good fit in a bin correspond to a value of $R$ close to 00, meaning that the simulation results are close to the observations. Inspecting fig. 7, we note a stronger deviation from the observations at the boundaries of the density plot, for instance it confirms too many pulsars are piling up close to the death line. The high values of $R$ at these boundaries originates from the weak number of pulsars counted in this bins leading to bad statistics and large variations in the proportions, a side effect of the density plot approach. The fact that the colors are very dark, shows that the proportion of pulsars in most of the bins are not good.
A Kolmogorov Smirnov (KS) test (Smirnov, 1948) was conducted on both $P$ and $\dot{P}$ which gave p-values substantially below 0.05 respectively when no death line is implemented. While we obtained p-values of 0.05 and another one well below 0.05 respectively for $P$ and $\dot{P}$ with a death line implementation. The null hypothesis that observations and simulations originate from the same distribution for the spin period can not be rejected (p-value $\geq$ 0.05) only when a death line is considered, but is clearly rejected for $\dot{P}$ in both case.
3.2 Simulating 1 millions pulsars
3.2.1 With the death line
The parameters used for these runs are shown in Table 5. This new set of simulations is intended to verify if generating pulsars younger than $4.1\times 10^{7}$ yr is sufficient, since we observe a small excess of very old pulsars. The death line is also taken into account.
Table 8: Number of pulsars detected with the death line implementation for 1 million pulsars simulated.
log($\dot{E}$) (W) $N_{tot}$ $N_{r}$ $N_{g}$ $N_{rg}$
¿31 2 0 2 0
¿28 124 7 73 44
Total 1694 1414 154 126
Figure 8: Distribution of the observed period taken from the ATNF catalogue, along with the simulations with the death line implementation for 1 million pulsars simulated.
Figure 9: Distribution of the observed period derivative taken from the ATNF catalogue, along with the simulations with the death line implementation for 1 million pulsars simulated.
Figure 10: $P-\dot{P}$ diagram of the simulated population, along with the observations, with the death line implementation for 1 million pulsars simulated.
Table 8 shows that a few less pulsars were detected in this simulation compared to the one with the death line and 10 millions pulsars (see Table 7). Moreover, inspecting the histograms for spin period and spin period derivative respectively in Fig. 8 and 9, we conclude that observations and simulations are very similar in both histograms. A KS test was conducted for both distributions $P$ and $\dot{P}$ and yielded p-values of $0.24$ and $0.07$ respectively. Therefore, it is impossible to reject the null hypothesis that the simulated distributions are similar to the observed distribution.
Figure 11: Density plot of the $P-\dot{P}$ diagram in comparison with observations for 1 million pulsars simulated.
Fig 10 shows the good agreement between simulations and observations. Furthermore, we can compare now Fig. 11 and Fig. 7. It is obvious that Fig. 11 is better, because globally the colors are lighter which signifies the proportion of pulsars in each 2D bin are almost alike. In addition, there is not a pile up effect anymore next to the death line, which corresponds better to the data.
Figure 12: Distance distribution to Earth for observed and simulated populations (1 million pulsars), including the death line implementation.
The distribution of the distance to Earth is shown in Fig. 12. It has a slightly better similarity to the observations than the work of Dirson et al. (2022). Nevertheless the improvement is marginal compared to the improvements of $P$ and $\dot{P}$ histograms. However, we noticed that removing the effect of dispersion and scattering of the ISM on the radio pulse profile, we would obtain very similar distribution of distances between observations and simulations. Therefore, a refinement in the description of the effect of the ISM on the pulse profile might be necessary.
Indeed, Fig. 13 confirms that the positions of the pulsars in the X-Y plane of the Galaxy are slightly more scattered than before, but the improvement is marginal.
Figure 13: Spatial distribution of the simulated pulsars, along with the observations, projected onto the Galactic plane for 1 million pulsars simulated.
Figure 14: Distribution of the observed latitude taken from the ATNF catalogue, along with the simulations with the death line implementation for 1 million pulsars simulated.
Finally fig. 14 compares the Galactic latitudes from the previous study (see Fig. 9 of Dirson et al. (2022)), now the simulated distribution resembles more the observations, especially around all the latitudes that are different from 0°. Therefore, the Galactic potential helped tracking the pulsar trajectories more realistically.
3.2.2 Without a death line
Because the simulation where 1 million pulsars are simulated do not allow to obtain pulsars older than $4.1\times 10^{7}$ yr, we undertook a thorough examination of the usefulness of a death line in modeling the canonical pulsar population. Upon conducting simulations with and without a death line (see Table 5 for the parameter values used in the simulations), it emerged that the resultant number of detected pulsars and the corresponding plots exhibited remarkable similarities. In addition, when we conduct the KS test on $P$ and $\dot{P}$ between 2 simulations with and without death line, we get p-values of 0.71 and 0.16 respectively, which confirms the similarity between the two simulations. Consequently, our attention turned towards exploring the impact of lower birth rates to ascertain if they provide a more congruous fit for both spin period and spin period derivative distributions.
Figure 15: $P-\dot{P}$ diagram of the simulated population, along with the observations, without death line implementation, birth rate of 1/70 yr-1 for 1 million pulsars simulated.
Figure 16: $P-\dot{P}$ diagram of the simulated population, along with the observations, without death line implementation, birth rate of 1/150 yr-1 for 1 million pulsars simulated.
Running simulations with a birth rate of $1/70$ yr-1 or a birth rate of $1/150$ yr-1 (as a consequence, increasing the oldest simulated pulsar), conspicuously revealed an insufficient number of detected pulsars in both Fig. 15 and Fig. 16. Intriguingly, a discernible paucity of pulsars congregating in the bottom right quadrant of the diagram, colloquially known as the pulsar graveyard, persisted even under these conditions, therefore regardless of the birth rates, thus casting doubt on the necessity of a death line in this configuration. Ultimately, our analysis underscores the pivotal importance of delineating the maximum age of simulated pulsars, as elucidated in Subsection 3.1, while the incorporation of a death line do not seem indispensable for attaining realistic results when simulating 1 million pulsars.
3.3 Spin-velocity misalignment effect
With the aim of observing the effect of the Galactic potential, we looked at the alignment of the rotation axis and the proper motion velocities of the pulsars. Results are taken from the simulation containing 1 million pulsars, see Table 5 for the parameters used. Noutsos et al. (2013) showed by studying a catalog of 58 pulsars that pulsars younger than 10 Myr keep their rotation axis and their velocities vector still aligned or anti-aligned, because they were aligned at birth according to Rankin (2007). Meanwhile, pulsars older than 10 Myr do not show this trend anymore. This progressive misalignment effect is due to the movement through the Galactic potential.
Figure 17: Distribution of the spin-velocity angles for all the detected pulsars in the simulation for 1 million pulsars simulated.
Figure 18: Distribution of the spin-velocity angles for the detected pulsars younger than 10 Myr (blue) and pulsars older than 10 Myr (red) in the simulation.
Fig. 17 demonstrates that most pulsars have their rotation axis aligned with their velocities, but for spin-velocity angles between 25° and 180° the numbers of pulsars is considerably lower than around 0° and 180°. On the one hand, for pulsars younger than 10 Myr, shown in blue in Fig. 18, only a small fraction of pulsars have their spin-velocity angle greater than 25° and below 150°. On the other hand, for pulsars older than 10 Myr, shown in red in Fig. 18, their spin-velocity angle becomes greater than 25° or below 150°.
Figure 19: The spin-velocity angles for the detected pulsars in function of the number of orbits made by the pulsars.
Furthermore, in Fig. 19, we can confirm that, the older the pulsar is, the better is its chance to orbit more in the Galaxy, allowing it to deviate its spin-velocity angle from alignment (or anti-alignment). In addition, even if the pulsar is not that old, if it made many orbits in the Galaxy, its spin-velocity angle will be between 25° and 150°, which confirms the fact that its the Galactic potential (and not only age) which makes this spin-velocity angle increase (or decrease if its is anti-aligned at birth). In the contrary, there are pulsars which are old which did not made many orbits in the Galaxy, as a consequence they have their spin-velocity angle near 0° or 180°. These plots clearly shows the influence of the Galactic potential making the pulsars lose their alignment between their rotation axis and their velocities through time because they evolve in the Galactic potential.
3.4 Log normal vs normal distribution for spin period at birth
Figure 20: Distribution of the observed period taken from the ATNF catalogue, along with the simulations with the death line implementation and the Gaussian distribution at birth for the spin period.
Table 9: Number of pulsars detected with the death line implementation and the Gaussian distribution at birth for the spin period.
log($\dot{E}$) (W) $N_{tot}$ $N_{r}$ $N_{g}$ $N_{rg}$
¿31 7 0 5 2
¿28 275 9 138 128
Total 1642 1188 223 231
In order to estimate the impact of a new distribution of spin period at birth, we compare the results with the previous distribution of spin period
$p(P_{0})=\frac{1}{\sigma_{p}\sqrt{2\pi}}e^{-(P_{0}-\bar{P})^{2}/(2\sigma_{p})^{2}}$ (34)
The simulation in this subsection uses $\bar{P}=60$ ms and $\sigma_{p}=10$ ms as parameters of equation (34). The Galactic potential and the death line were both taken into account to compare the influence of the initial spin period distribution and 1 million pulsars were simulated, see Table 5 for the parameters used.
Table 9 highlights the fact that much more pulsars are detected above $\dot{E}>10^{28}$ W compared to the simulation with the Gaussian in log space for the spin period, see Table 7. It confirms the suspicion that the distribution of initial spin period influences the number of pulsars at higher spin down luminosity.
The spin period distribution in Fig. 20 shows some simulated pulsars with a low spin period, in excess compared to the observations. This artifact disappears when the Gaussian distribution in log space for the spin period is used. The log-normal distribution enables to remove those pulsars in excess and make the distribution of the spin period better fit the observations. The KS test was also conducted on both $P$ and $\dot{P}$ and we find p-values, once again, considerably below 0.05 for both quantity, therefore rejecting the null hypothesis. We conclude that the initial distribution for the spin period found by Igoshev et al. (2022) allows to better reproduce the canonical pulsars population.
3.5 Gamma-ray detection
Let us now focus on the $\gamma$-ray band of the pulsar population. First, we note that the total number of pulsars detected simultaneously in gamma-ray and in radio is significantly larger in our simulation compared to the observations, when using the Fermi/LAT instrument sensitivity. However, the simulated gamma-ray pulsars are in the expected area of Fig. 21, meaning the striped wind model allow to correctly reproduce these pulsars. The explanation of this higher number of detection would be that there are many unidentified sources in the Fourth Fermi-LAT Catalog of Gamma-ray sources (4FGL). Since many unidentified Galactic sources are listed in these catalogs, many gamma-ray pulsars might have been detected without being identified as such yet. Furthermore, we can compare our results with the PPS work of Pierbattista et al. (2012), where they simulated the young gamma-ray pulsars population with 4 different models : Outer Gap (OG), Slot Gap (SG), Polar Cap (PC) and Outer Polar Cap (OPC) models. In summary their best model was the OPC model, which shows almost identical results as ours (the $P-\dot{P}$ shape and the number of detection), except the fact that we detect gamma-ray pulsars a little further away from Earth.
Figure 21: $P-\dot{P}$ diagram of the gamma-only pulsars for both the simulations and the observations, with 1 million pulsars simulated.
What if we run another simulation (with the same parameters as in Table 5), by only changing the threshold of detection for gamma-ray ? We did this in order to see how many more pulsars we would detect in gamma-ray and where they would be in the $P-\dot{P}$ diagram, and also to deal with the problematic of the GeV excess in the Galactic centre. Therefore we increased $F_{\rm min}$ by 10 : if the galactic latitudes of the pulsar is ¡ 2°, then $F_{\rm min}=0.4\times 10^{-15}$ W.m-2, and concerning blind searches $F_{\rm min}=1.6\times 10^{-15}$ W.m-2.
Figure 22: Spatial distribution of the simulated gamma-ray detected pulsars, projected onto the Galactic plane for 1 million pulsars simulated, with a lower $F_{min}$.
The gamma-ray detected pulsar positions in the X-Y plane of the Galaxy are shown in Fig. 22. In this simulation, 1832 gamma-ray pulsars are detected, among them only 168, in red in Fig. 22, would have been detected if the same sensitivity as before was kept. Therefore we detect 11 times more gamma-ray pulsars compared to the number found with the previous sensitivity. Furthermore, the other 1664 pulsars, in green in Fig. 22, detected with this increased sensitivity highlight several features. We obviously detect more pulsars close to us, and also many more in the centre of the Galaxy. This result indicates that the GeV excess in the Galactic centre could be linked to pulsars not yet identified, while usually the hypothesis is that the GeV excess corresponds to self-annihilating dark matter particles (Hooper, 2022). In addition, millisecond pulsars, not modeled in this work, could however also contribute to this GeV excess.
4 Discussion
For the initial spin period distribution Igoshev et al. (2022) suggested values for the mean period $P_{\text{mean}}$ in the range from approximately $57$ ms to $129$ ms and for the spread $\sigma_{p}$ the range from $0.45$ to $0.65$. The values $P_{\text{mean}}=129$ ms and $\sigma_{p}=0.45$ are best educated guesses. However it is obvious that further investigations could benefit from an automated method to find the best set of values. Hence, we are currently working on an optimization algorithm to find the optimal initial parameters for the population synthesis, not only for the spin period, but for the whole parameter space presented in Table 5, except for $\alpha_{d}$ and $k_{\tau_{d}}$. Our optimization algorithm will help to constrain the properties of the Milky Way pulsar population by constraining parameters like the birth rate which is supposed to be between $1/33$ yr-1 and $1/150$ yr-1, see Faucher-Giguère & Kaspi (2006), Gullon et al. (2014) or Johnston & Karastergiou (2017).
We also added the effect of the ISM on the radio pulse profile, rendering more difficult to observe pulsars farther away than $10$ kpc. Furthermore, because of this plasma propagation effect fewer radio pulsars are detected compared to the observations. If the simulations discard the ISM effect, much more pulsars are detected at distances larger than $10$ kpc and with higher $\dot{E}$ as well, improving the match between the simulation and the observation. Perhaps, the effect of the ISM is overestimated when computing the scattering and dispersion effect converted in equation (28a) with $\tau_{scat}$ and $\tau_{DM}$ is too important and should have less influence on the observed width pulse profile, therefore with formulas that would give lower $\tau_{DM}$ and $\tau_{scat}$, we might get more realistic results.
In the main part of the paper, we abstain from showing outcomes stemming from simulations featuring a constant magnetic field. However we ran such simulations that clearly highlight the inferior performance of a simulation without magnetic field evolution as can be checked in appendix B. Moreover appendix C explores different parameter spaces for the PPS in order to meticulously scrutinizing whether the parameters showcased in the main part of the paper optimally encapsulate observational data. This demonstrates that even if a substantial number of parameters are required for the PPS, they can be meaningfully constrained by such studies with reasonable accuracy.
Furthermore, in Subsection 3.5 we explore the results when the sensitivity in gamma-ray is increased by a factor of 10. However we did not show what would happen if we tried to make a prediction for the radio detection by increasing the sensitivity by a factor of 50 to match approximately the properties of SKA. A simulation was run with this increased sensitivity and 52276 pulsars were detected (when simulating 1 million pulsars, with a death line), therefore approximately 31 times more than with the sensitivity of Parkes and Arecibo telescopes.
5 Summary
Without the implementation of the Galactic potential, in the PPS of Dirson et al. (2022), the pulsars were moving in straight lines along one direction chosen randomly for each pulsar. This evolution of the proper motion was not realistic, even though the evolution model had already replicated a significant portion of the canonical pulsar population. However all the new improvements put into this PPS, meaning : the Galactic potential, the effect of the ISM on the radio pulse profiles and the fact that the decay timescale of the magnetic field is taken randomly within a certain range, allow a more accurate reproduction of the population in the $P-\dot{P}$ diagram.
The first results where we emulate 10 millions pulsars showed an excess of old pulsars detected in the simulation. In the $P-\dot{P}$ diagram we saw that this excess was associated to pulsars which lie below the death line, therefore, not creating pairs anymore. With this finding we realized that it was imperative to implement the death line. The results obtained with the death line were then in better agreement with observations when plotting the most relevant distributions such as period and period derivative.
Nonetheless, simulating 1 million pulsars allowed to show that with or without death line the results were alike and therefore, in the case where the oldest pulsar simulated is $4.1\times 10^{7}$ yr, a death line is not really necessary to explain the observations. In this paper we showed two approach to reproduce the observed canonical pulsars population, however when we consider the approach where the oldest pulsar simulated is only $4.1\times 10^{7}$ yr we state that no older pulsars might exist in the Galaxy, while the Milky Way is very much older than this. There is apparently no reason to not simulate pulsars that are older than this, and if there is one we do not know it (they become harder to detect because of their wide pulse profile with their age, but some are still detectable). If this situation is the reality then it means two things : the age of the observed pulsars are overestimated (because we observe pulsars older than $4.1\times 10^{7}$ yr in the canonical population, even if they are not numerous) and the death line is not necessary in pulsar population synthesis work to explain what we observe. When the simulation is ran with 10 millions pulsars, the death line becomes a necessity. However this simulation gives results less similar to the observations, because of a few simulated data which do not match any observational data.
It was also shown that the old pulsars with age larger than 10 Myr have a tendency to lose their alignment between rotation axis and velocity, a direct effect of the Galactic potential as shown by the observations of Noutsos et al. (2013), but it is the first time it is done in a simulation.
The striped wind model seems to reproduce the gamma-ray population of canonical pulsars very well, however in the simulation a higher number of these pulsars are detected compared to reality. It probably indicates that a part of unidentified sources from the 4FGL catalog are pulsars. Testing an increased sensitivity for the detection of gamma-ray pulsars lead us to think that a part of the origin of the GeV excess in the Galactic centre could be young gamma-ray pulsars, and when the millisecond pulsars will be modeled in our simulation we would also be able to check that these pulsars might be responsible for this GeV excess. The usual picture is to suppose that only millisecond pulsars are responsible, while young gamma-ray pulsars could also play a role.
In summary, the best parameters found for this pulsar population synthesis are for $P_{\text{mean}}=129$ ms, $\sigma_{p}=0.45$ for the spin period distribution at birth, $B_{\text{mean}}=2.75\times 10^{8}$ T, $\sigma_{b}=0.5$ for the magnetic field distribution at birth and for a birth rate of 1/41 yr-1. Ultimately, this work provides a simulated population of pulsars within the Milky Way more similar to the observations compared to Johnston et al. (2020) and Dirson et al. (2022), the two most recent population synthesis considering both radio and gamma-ray emission. Moreover, such investigations will be useful to make more predictions for future radio surveys such as SKA, or NenuFAR.
The PPS method used in this study could be applied to model populations of pulsars in the Milky Way other than the canonical ones. The magnetars, for instance, would need : another distribution at birth for their magnetic field to be modeled correctly, taking into account the emission in X-rays and having a faster decay for the magnetic field. The magnetars are the pulsars just above the canonical pulsars in the $P-\dot{P}$ diagram, at the top right of the diagram. A more complex population is the recycled or millisecond pulsars, that were ”brought back to life” by accreting matter from a companion star. This accretion mechanism spins up the neutron stars moving them from the graveyard of pulsars to the bottom left of the $P-\dot{P}$ diagram, crossing again the death line in the opposite sense and allowing them to create electron-positron pair cascades again. However, they are more complex to model because this evolution scenario should take into account the accretion process and spin-up. We are also currently working on both populations to reproduce them in a PPS based on this work.
Acknowledgements.
This work has been supported by the grant ANR-20-CE31-0010. We acknowledge the High Performance Computing Centre of the University of Strasbourg for supporting this work by providing scientific support and access to computing resources. D.M. acknowledges the support of the Department of Atomic Energy, Government of India, under project No. 12-R&DTFR-5.02-0700.
Appendix A The integration scheme used
A.1 The scheme
One of the new aspect in this work is to solve the equation of motion (12) in order to follow more realistically the motion of our sample of pulsars. To achieve this objective we implemented a fourth order integration scheme called Position Extended Forest Ruth-Like (PEFRL). This algorithm was chosen because of its high order allowing a high precision on the trajectory of the pulsar. In addition, it is a symplectic algorithm, meaning that it conserves the total energy of the pulsar (kinetic plus potential), minimizing the error even if the integration is done for a long time with any orbits. It works in the same way as a Verlet scheme of second order, the velocities and coordinates are shifted separately but in the end they get synchronized. The different steps of the algorithm are as follows
${r}^{n+\frac{1}{5}}$ =${r}^{n}+\xi h\@vec{v}^{n}$ (35a)
${v}^{n+\frac{1}{4}}$ =${v}^{n}-(1-2\lambda)\frac{h}{2}\@vec{\nabla}\Phi(\@vec{r}^{n+\frac{1}{5}})$ (35b)
${r}^{n+\frac{2}{5}}$ =${r}^{n+\frac{1}{5}}+\chi h\@vec{v}^{n+\frac{1}{4}}$ (35c)
${v}^{n+\frac{2}{4}}$ =${v}^{n+\frac{1}{4}}-\lambda h\@vec{\nabla}\Phi(\@vec{r}^{n+\frac{2}{5}})$ (35d)
${r}^{n+\frac{3}{5}}$ =${r}^{n+\frac{2}{5}}+(1-2(\chi+\xi))h\@vec{v}^{n+\frac{2}{4}}$ (35e)
${v}^{n+\frac{3}{4}}$ =${v}^{n+\frac{2}{4}}-\lambda h\@vec{\nabla}\Phi(\@vec{r}^{n+\frac{3}{5}})$ (35f)
${r}^{n+\frac{4}{5}}$ =${r}^{n+\frac{3}{5}}+\chi h\@vec{v}^{n+\frac{3}{4}}$ (35g)
${v}^{n+1}$ =${v}^{n+\frac{3}{4}}-(1-2\lambda)\frac{h}{2}\@vec{\nabla}\Phi(\@vec{r}^{n+\frac{4}{5}})$ (35h)
${r}^{n+1}$ =${r}^{n+\frac{4}{5}}+\xi h\@vec{v}^{n+1}$ (35i)
here ${r}$ is the position vector, ${v}$ is the velocity vector, $h$ is the time step, ${\nabla}\Phi$ is the gradient of the potential and $n$ is the time at which we compute the position or the velocity. $\lambda$, $\xi$ and $\chi$ are constants that are tabulated in Table 10, see for instance Peter Young (2014).
Table 10: Value of the constants used for the PEFRL scheme.
Parameters Values
$\lambda$0.212341831062605
$\xi$ 0.178617895844809
$\chi$0.06626458266981849
A.2 Precision of the PEFRL algorithm
Before implementing the gravitational potential and the integration scheme in the whole synthesis, the precision of the trajectory of a single pulsar in the Milky Way was checked. The initial conditions for the pulsar are the following : $x=5.1$ kpc, $y=-1.9$ kpc, $z=0.05$ kpc, $v_{x}=-47.8$ km/s $v_{y}=97.0$ km/s, and $v_{z}$ = 22.6 km/s. Those initial conditions were chosen randomly, the only constrain being to have the pulsar in the Galaxy with realistic coordinates on the one hand and with realistic speeds for a pulsar on the other hand. Furthermore, the total integration time is 1.25 Gyr and the time step used is $10^{5}$ yr. In order to check the precision of the scheme, the total energy of the pulsar, sum of the kinetic plus potential energy, that should be conserved is computed as
$E_{tot}=\frac{v_{x}^{2}}{2}+\frac{v_{y}^{2}}{2}+\frac{v_{z}^{2}}{2}+\Phi_{tot}.$ (36)
The trajectory obtained after the integration with the PEFRL algorithm is shown in Fig. 23. This trajectory shows that the pulsar, with these initial conditions, remains bound to the Galaxy. The pulsar here takes approximately 201.3 Myr to complete an orbit in the Galaxy. The time step chosen is significantly smaller than the orbital period ($h/P_{\rm orb}=0.0005$), and we obtain approximately 6 orbits by integrating the trajectory during 1.25 Gyr. Bajkova & Bobylev (2021) showed several trajectories in their paper, the one obtained here seems coherent compared to the different trajectories that they got.
Figure 23: Trajectory of a single pulsar in the Galactocentric frame.
Fig. 24 shows the relative error on the total energy. It is indeed almost constant along the whole integration of the trajectory. The maximum relative error on the energy is of the order of $10^{-9}$, a very good accuracy for this trajectory. Globally, when a single pulsar has its trajectory integrated in the PPS, the relative error is at maximum 0.001 and the minimum seen was $10^{-14}$.
Figure 24: Relative error on the energy of a single pulsar in the Galactocentric frame.
Appendix B Results for a constant magnetic field
Depending on the PPS, the magnetic field will be considered constant (Johnston & Karastergiou 2017; Faucher-Giguère & Kaspi 2006) or decaying (Gullon et al. 2014; Dirson et al. 2022; Jawor & Tauris 2022). The goal is here is to show the results with a constant magnetic field in order to see which model is better between the decaying magnetic field and the constant magnetic field.
Hence, this simulation was run with 1 million pulsar simulated, without a death line and without a decaying of the magnetic field. The KS test gave p-values of $6.3\times 10^{-22}$ and $9.8\times 10^{-77}$, therefore the null hypothesis is rejected for both $P$ and $\dot{P}$. The parameters used are in Table 5. It is less good than with a decaying magnetic field. As can be seen in Fig. 25, too many pulsars from the simulation appear on the top right of the diagram and almost none at the bottom. In Table 11, we note a total of pulsars which is smaller from what we found for all of our previous results in Sect. 3.
Figure 25: $P-\dot{P}$ diagram of the simulated population, along with the observations, with the death line implementation and a constant magnetic field for each pulsar.
Table 11: Number of pulsars detected for the simulation with constant magnetic field for each pulsar and the death line implementation.
log($\dot{E}$) (W) $N_{tot}$ $N_{r}$ $N_{g}$ $N_{rg}$
¿31 15 0 12 3
¿28 168 4 94 70
Total 1356 1034 162 160
Appendix C Influence of the different parameters
In order to show that the parameters chosen in the main part of this paper are meticulously picked and that the uncertainties brought by the large number of parameters is manageable, several simulations were run with different values for the main parameters of the initial distributions, showing a deviation from the best values gives bad results. The simulations were run with in the same conditions as Subsection 3.2, with only one parameter changing between the different sets of simulation.
Table 12: Influence of the birth rate.
Birth rate (in yr-1) $N_{tot}$ $N_{r}$ $N_{g}$ $N_{rg}$
1/30 2388 1964 204 220
1/80 963 799 79 85
1/120 696 597 50 49
Table 13: Influence of the mean initial period.
$P_{\rm mean}$ (in ms) $N_{tot}$ $N_{r}$ $N_{g}$ $N_{rg}$
57 1499 1134 193 172
70 1623 1253 215 155
100 1652 1348 151 153
Table 14: Influence of the mean initial magnetic field.
$B_{mean}$ (in T) $N_{tot}$ $N_{r}$ $N_{g}$ $N_{rg}$
$10^{7}$ 8407 6943 406 1058
$5\times 10^{7}$ 4447 3646 306 495
$5\times 10^{8}$ $1$ 933 109 69
The $P-\dot{P}$ diagram of each simulation is not shown here because almost every simulation finds a too high (or too low) number of pulsars, therefore the $P-\dot{P}$ diagrams are showing large differences between observations and simulations. Nonetheless, the number of pulsars detected for two of the simulations with different $P_{\rm mean}$ (with $P_{\rm mean}=70$ ms and $P_{\rm mean}=100$ ms) gives number closer to reality. However their p-values for $P$ and $\dot{P}$ when using the KS test were significantly below $0.05$, which allowed us to reject the null hypothesis for the distributions of $P$ and $\dot{P}$, for the results shown in Table 12, 13 and 14.
Moreover in order to get an idea on a good range of values for the parameters, the d-value obtained thanks to the KS test are shown in Table 15, 16, 17 and 18. For each parameter, 10 simulations were made in order to get a mean and a standard deviation for the d-value associated. These simulations helped finding the best values for the parameters by considering the parameter is well constrained when its d-values for both distribution of $P$ and $\dot{P}$ are below, or really close to 0.05 (meaning 5 % similarity between the observations and simulation distribution). Even though it is not an optimisation (and therefore the results can be improved on that point), it allows us to conclude that $P_{\rm mean}$ should be between 120 ms and 135 ms (with 129 ms as best value). $B_{\rm mean}$ should be between $2.25\times 10^{8}$ T and $2.8\times 10^{8}$ T (with $2.75\times 10^{8}$ T as best value). The birth rate should be between 1/37 yr-1 and 1/45 yr-1 (with 1/41 yr-1 as best value). Finally, $\sigma_{b}$ is well constrained between 0.49 and 0.51 (with 0.5 as best value). Concerning $\sigma_{p}$, it always gives d-values between 0.04 and 0.057 when $\sigma_{p}$ is below 0.6, therefore the value 0.45 was chosen as suggested by Igoshev et al. (2022).
Table 15: D-value for different $P_{\rm mean}$.
$P_{\rm mean}$ (in ms) d-value ($\dot{P}$) d-value ($P$)
90 $0.072\pm 0.0073$ $0.0577\pm 0.0080$
100 $0.0651\pm 0.00692$ $0.0503\pm 0.00636$
110 $0.0652\pm 0.00688$ $0.0381\pm 0.0074$
120 $0.0554\pm 0.00714$ $0.0302\pm 0.00468$
129 $0.0557\pm 0.00740$ $0.0284\pm 0.00583$
135 $0.0544\pm 0.00768$ $0.0329\pm 0.00767$
150 $0.0572\pm 0.00779$ $0.0388\pm 0.00935$
Table 16: D-value for different $B_{\rm mean}$.
$B_{\rm mean}$ (in T) d-value ($\dot{P}$) d-value ($P$)
 $1\times 10^{8}$ $0.113\pm 0.00823$ $0.203\pm 0.00675$
$2\times 10^{8}$ $0.0418\pm 0.00459$ $0.0658\pm 0.00705$
$2.25\times 10^{8}$ $0.0456\pm 0.00455$ $0.0455\pm 0.0061$
$2.75\times 10^{8}$ $0.0557\pm 0.00740$ $0.0284\pm 0.00583$
$2.8\times 10^{8}$ $0.0576\pm 0.00907$ $0.0298\pm 0.00728$
$3\times 10^{8}$ $0.0648\pm 0.01092$ $0.0426\pm 0.00775$
$3.25\times 10^{8}$ $0.0761\pm 0.00930$ $0.0480\pm 0.00997$
Table 17: D-value for different $\sigma_{b}$.
$\sigma_{b}$ d-value ($\dot{P}$) d-value ($P$)
0.45 $0.0767\pm 0.00668$ $0.0513\pm 0.01258$
0.48 $0.0621\pm 0.00880$ $0.0385\pm 0.01234$
0.49 $0.0646\pm 0.00974$ $0.0304\pm 0.00501$
0.5 $0.0557\pm 0.00740$ $0.0284\pm 0.00583$
0.51 $0.056\pm 0.00735$ $0.0306\pm 0.00701$
0.52 $0.0547\pm 0.00786$ $0.0358\pm 0.00579$
Table 18: D-value for different birth rate (BR).
BR (in yr -1) d-value ($\dot{P}$) d-value ($P$)
1/33 $0.0611\pm 0.01025$ $0.0313\pm 0.00636$
1/37 $0.0565\pm 0.00851$ $0.0295\pm 0.00645$
1/41 $0.0557\pm 0.00740$ $0.0284\pm 0.00583$
1/45 $0.0585\pm 0.00510$ $0.0302\pm 0.00565$
1/70 $0.0691\pm 0.00830$ $0.0342\pm 0.01165$
1/150 $0.1165\pm 0.01564$ $0.0408\pm 0.00671$

### v2

@@
 Pulsars are highly magnetized rotating neutron stars, emitting in a broad electromagnetic energy range.
-These objects have been discovered more than 50 years ago and are astrophysical laboratories for studying physics at extreme conditions.
-Reproducing the observed population of pulsars refines our understanding of their formation, evolution as well as radiation processes and geometry.
+These objects have been discovered more than 55 years ago and are astrophysical laboratories for studying physics at extreme conditions.
+Reproducing the observed pulsars population refines our understanding of their formation and evolution scenarios as well as their radiation processes and geometry.
 Aims.
-In this paper, we improve our previous pulsar population synthesis by studying the impact of the Galactic gravitational potential and of the death line putting emphasis on the $\gamma$-ray pulsar population.
-In order to elucidate the necessity of a death line, refined initial distributions for spin periods and radial positions at birth were implemented, elevating the sophistication of our previous simulations to the most recent state-of-the-art.
+In this paper, we improve our previous population synthesis by focusing on both the radio and $\gamma$-ray pulsar populations, investigating the impact of the Galactic gravitational potential and of the radio emission death line.
+In order to elucidate the necessity of a death line, refined initial distributions of spin period and spacial position at birth were implemented, elevating the sophistication of our simulations to the most recent state-of-the-art.
 Methods.
-The individual motion of pulsars in the Galactic potential is evolved by a fourth order symplectic integration scheme.
-Our pulsar population synthesis takes into account the secular evolution of a force-free magnetosphere and the magnetic field decay.
+The motion of each individual pulsar is tracked in the Galactic potential by a fourth order symplectic integration scheme.
+Our pulsar population synthesis takes into account the secular evolution of the force-free magnetosphere and magnetic field decay simultaneously and self-consistently.
 Each pulsar is evolved from its birth up to the present time.
-The radio and $\gamma$-ray emissions are modeled respectively by the polar cap geometry and the striped wind model.
+The radio and $\gamma$-ray emission locations are modelled respectively by the polar cap geometry and the striped wind model.
 Results.
-When simulating a cohort of ten million pulsars akin to previous endeavors, the integration of a death line proves essential for a better agreement with observational trends.
-Yet, discrepancies emerges when modeling a reduced ensemble of only one million pulsars, endowed with a higher birth rate.
-This divergence yields a remarkably refined $P-\dot{P}$ diagram, conspicuously with or without any delineating death line.
-Intriguingly, this observation suggests an overestimation of the ages of detected pulsars in existent data sets, thereby challenging the necessity of death lines in pulsar population synthesis.
-Validation through Kolmogorov-Smirnov testing underlines the statistical congruence between the spin period and spin period derivative distributions of both observations and simulations.
-Moreover simulations with a higher $\gamma$-ray instrumental sensitivity show that the GeV excess in the Galactic centre could be of pulsar origin.
+By simulating ten million pulsars we found that including a death line better reproduces the observational trend.
+However, when simulating one million pulsars, we obtain an even more realistic $P-\dot{P}$ diagram, whether or not a death line is included.
+This suggests that the ages of the detected pulsars might be overestimated, therefore questioning the real need for a death line in pulsar population studies.
+Kolmogorov-Smirnov tests confirm the statistical similarity between the observed and simulated $P-\dot{P}$ diagram.
+Additionally, simulations with increased $\gamma$-ray telescope sensitivities hint to a significant contribution of $\gamma$-ray pulsars to the GeV excess in the Galactic centre.
 Key Words.:
 pulsars: general - Gravitation - radio continuum: stars - Gamma rays: stars - methods: statistical
+[…]
 1 Introduction
-[…]
-Pulsars, discovered by Jocelyn Bell Burnell in 1967 (Hewish et al., 1968) are rotating neutron stars born in a core-collapsed supernova.
-Pulsars are highly magnetized and surrounded by a plasma-filled magnetosphere emitting regular pulses of radiation at their spin frequency.
-Due to the magneto-dipole losses, they lose energy as their spin period increases.
-Thanks to radio surveys, performed for instance by the Parkes and Arecibo radio-telescopes, nearly 3000 pulsars have been discovered (Manchester et al., 2005) (see Table 1, with the data on the canonical pulsars).
-Although pulsars were first observed in radio, it was found later that they are also bright in X-rays, optical and gamma-rays.
-The Large Area Telescope (LAT) on board of the Fermi satellite, has discovered dozens of radio-quiet gamma-ray pulsars as well as millisecond pulsars (MSPs).
-Since its launch, the LAT has detected about 300 pulsars (Abdo et al., 2013; Smith et al., 2019, 2023).
-The LAT has unveiled a new perspective, expanding our understanding about pulsars by presenting a growing collection of neutron stars that are detectable through their high energy emission.
-Moreover, the distinct emission process in high-energy and its unique beaming properties offer an alternative viewpoint on the global pulsar population.
-Table 1: Number of known pulsars with spindown luminosity $\dot{E}$ above and below $10^{28}$ W, and above $10^{31}$ W.
-The quantities $N_{r}$, $N_{g}$, and $N_{rg}$ are the number of radio-only, gamma only, and radio-loud gamma-ray pulsars, respectively.
-It should be noted that we excluded the binary pulsars as well as pulsars with $P<20$ ms.
-The data have been taken from the ATNF catalogue and from https://confluence.slac.stanford.edu/display/ GLAMCOG/Public+List+of+LAT-Detected+Gamma-Ray+Pulsars
-log($\dot{E}$) (W) $N_{tot}$ $N_{r}$ $N_{g}$ $N_{rg}$
-¿31 2 0 0 2
-¿28 197 101 35 61
-Total 2665 2553 63 84
-With the advent of future surveys such as the Square Kilometre Array (SKA  [^1] https://www.skao.int/en/science-users/118/ska-telescope-specifications) or the Cherenkov Telescope Array (CTA  [^2] https://www.cta-observatory.org/about/how-ctao-works/), much more pulsars are expected to be detected.
+Pulsars, discovered by Jocelyn Bell Burnell in 1967 (Hewish et al., 1968) are rotating neutron stars born in a core-collapse supernova.
+They are highly magnetized and surrounded by a plasma-filled magnetosphere emitting regular pulses of radiation at their spin frequency.
+Due to the magneto-dipole losses, they lose rotational kinetic energy and their spin period increases.
+Extensive radio surveys, performed for instance by the Parkes and Arecibo radio-telescopes, so far 3700 pulsars have been discovered (Manchester et al.
+2005, see Table 1 listing the canonical pulsars).
+Although pulsars were first observed in radio, they were later found to be also bright in X-rays, optical and $\gamma$-rays.
+The Large Area Telescope (LAT) on board of the Fermi satellite, has discovered dozens of radio-quiet $\gamma$-ray pulsars as well as millisecond pulsars (MSPs).
+Since its launch, the LAT has detected about 300 pulsars (Abdo et al., 2013; Smith et al., 2019, 2023) and unveiled a new perspective, expanding our understanding about pulsar emission mechanisms by increasing the sample of neutron stars detectable through their high energy emission.
+Moreover, the distinct radiation process involved in high-energy photon production and its unique beaming properties offer an alternative viewpoint on the global pulsar population.
+Table 1: Number of known canonical pulsars, with spindown luminosity $\dot{E}$ above and below $10^{28}$ W, and above $10^{31}$ W.
+The quantities $N_{\rm r}$, $N_{\rm g}$, and $N_{\rm rg}$ are the number of radio-only, gamma only, and radio-loud $\gamma$-ray pulsars, respectively.
+It should be noted that we excluded the binary pulsars, the pulsars in the Magellanic clouds and in globular clusters.
+The data have been taken from the ATNF catalogue from https://www.atnf.csiro.au/research/pulsar/psrcat/
+log($\dot{E}$) (W) $N_{\rm tot}$ $N_{\rm r}$ $N_{\rm g}$ $N_{\rm rg}$
+¿31 4 0 2 2
+¿28 150 40 44 66
+Total 2193 2020 80 93
+With the advent of future surveys such as the Square Kilometre Array (SKA  [^1] https://www.skao.int/en/science-users/118/ska-telescope-specifications) or the Cherenkov Telescope Array (CTA  [^2] https://www.cta-observatory.org/about/how-ctao-works/), much more pulsars will be detected in both these wavelengths.
 Those instruments are scheduled to start collecting data by the end of the decade.
-SKA is going to be 50 times more sensitive and is predicted to survey the sky 10 000 times faster than any existing imaging radio telescope (McMullin et al., 2020).
-CTA is aimed at detecting gamma rays in the energy range from a few tens of GeV up to hundreds of TeV (Hofmann & Zanin, 2023) while Fermi/LAT operates in the energy range from 20 MeV to several hundred GeV  [^3] https://fermi.gsfc.nasa.gov/science/instruments/table1-1.html.
-Pulsar population synthesis (PPS) is a powerful tool aiming at predicting the discovery rate of new pulsars as well as better understanding their emission processes.
-PPS studies also constrains the overall properties of pulsars.
-Pulsars are generated from their birth and evolved up to the present.
-Once the set of pulsars has been generated we check whether they fulfill some prescribed detection criteria.
-The detectability of an individual pulsar in a given energy band (radio or $\gamma$) depends on whether the associated emission beam intersects our line of sight and on the sensitivity of the corresponding instrument.
-Most of PPS studies assume that neutron stars are rotating in vacuum (Faucher-Giguère & Kaspi, 2006; Popov et al., 2010; Johnston et al., 2020), or take only radio or gamma-ray emission into account (Watters & Romani, 2011; Gullon et al., 2014), or assume a constant magnetic field (Gonthier et al., 2002; Faucher-Giguère & Kaspi, 2006; Johnston & Karastergiou, 2017).
-For the first time in a PPS study, Dirson et al.
-(2022) have modeled the population of both the gamma-ray and radio pulsars by taking into account the force-free magnetosphere model in conjunction with the magnetic field decay.
+SKA is going to be 50 times more sensitive than current telescopes and will survey the sky 10,000 times faster than any existing imaging radio telescope (McMullin et al., 2020).
+CTA is aimed at detecting $\gamma$-rays in the energy range from a few tens of GeV up to hundreds of TeV (Hofmann & Zanin, 2023) while Fermi/LAT operates in the energy range from 20 MeV to several hundred GeV  [^3] https://fermi.gsfc.nasa.gov/science/instruments/table1-1.html and H.E.S.S operates in the energy range from 0.03 to 100 TeV  [^4] https://www.mpi-hd.mpg.de/HESS/pages/about/.
+Pulsar population synthesis (PPS) is a powerful tool to predict the discovery rate of new pulsars, to better understand their emission processes and to constrain their overall properties.
+In a PPS, pulsars are generated from their birth and evolved up to the present time.
+Once the sample of pulsars is generated, detection criteria are applied to check whether they can be detected by current and future telescopes.
+The detectability of an individual pulsar in a given energy band, radio or $\gamma$-ray, depends on whether the associated emission beam intersects our line of sight and on the sensitivity of the corresponding instrument.
+Most PPS studies assume that neutron stars are rotating in vacuum (Faucher-Giguère & Kaspi, 2006; Popov et al., 2010; Johnston et al., 2020) or take only radio or $\gamma$-ray emission into account separately (Watters & Romani, 2011; Gullon et al., 2014), or assume a constant magnetic field during the evolution process (Gonthier et al., 2002; Faucher-Giguère & Kaspi, 2006; Johnston & Karastergiou, 2017).
+Dirson et al.
+(2022) have modelled for the first time the population of both the $\gamma$-ray and radio pulsars, taking into account the state-of-the-art force-free magnetosphere model in conjunction with a prescription for the magnetic field decay.
 The neutron star radiation is produced by relativistic charged particles flowing within their magnetosphere.
-Therefore it is necessary to take the plasma back reaction into account for the pulsar period and dipole magnetic inclination angle evolution.Whereas the gamma-ray emission site from pulsars is still under debate, different locations have been suggested.
-These are for instance the polar cap region (Sturrock, 1971; Ruderman & Sutherland, 1975; Daugherty & Harding, 1982, 1996) at low altitudes, the slot gap along the last open magnetic field lines (Arons, 1983; Muslimov & Harding, 2004; Harding et al., 2008; Harding & Muslimov, 2011), or the outer gaps at high altitudes in the outer magnetosphere (Cheng et al., 1986; Hirotani, 2008; Takata et al., 2011).
-The striped wind model, which has for emission site the pulsar wind outside the light cylinder, (Pétri, 2009, 2011) is used for the first time by Dirson et al.
+Therefore the plasma back reaction must be taken into account in the evolution of the pulsar period and magnetic inclination angle.
+Whereas the location of the $\gamma$-ray emission site is still under debate, different regions have been suggested.
+These are for instance the polar cap region (Sturrock, 1971; Ruderman & Sutherland, 1975; Daugherty & Harding, 1982, 1996) at low altitudes, the slot gap along the last open magnetic field lines (Arons, 1983; Muslimov & Harding, 2004; Harding et al., 2008; Harding & Muslimov, 2011), or the outer gaps at high altitudes within the light-cylinder, in the outer magnetosphere (Cheng et al., 1986; Hirotani, 2008; Takata et al., 2011).
+Emission is also possible outside the light cylinder, in the so-called striped wind (Pétri, 2009, 2011) and was used for the first time by Dirson et al.
 (2022) in a PPS study.
 The radio emission is described as usual by the polar cap model.
-In this work, and the work of Dirson et al.
+In this work as in Dirson et al.
 (2022) only the canonical pulsar population in our Galaxy is reproduced.
-The canonical pulsar population is defined by isolated pulsars which are not magnetars, and with a spin period greater than 20 ms.
-While the study yielded satisfactory results when simulating a population closely resembling the observed pulsar population in the $P-\dot{P}$ diagram, there was room for improvement regarding the pulsar motion.
-Specifically, the study is limited by the assumption that pulsars move in straight lines at a constant speed from their birth to the present time.
-A noticeable discrepancy arises when comparing the observed position of pulsars to the simulation expectations, revealing a significant disparity in their spatial distributions.
-Since the results obtained by Dirson et al.
-(2022) did not require the modeling of a death line, it was not included in their PPS.
-The death line separates in the $P-\dot{P}$ diagram the neutron stars that generate particles through pair cascading from those which cannot radiate anymore and are hence called dead pulsars.
-This contrasts with the PPS performed by Graber et al.
-(2023) who did not require a death line.
-Therefore in this paper we examine the necessity of a death line to explain the observed canonical pulsars population.
-This paper improves the PPS of Dirson et al.
-(2022), questions the necessity of a death line in PPS work and discusses in depth the $\gamma$-ray pulsar population.
+The canonical pulsar population is defined by isolated pulsars which are neither magnetars nor millisecond pulsars.
+While the study by Dirson et al.
+(2022) yielded satisfactory results when simulating a population closely resembling the observed pulsar population in the $P-\dot{P}$ diagram, there was room for improvement regarding the pulsar trajectories within the Milky Way.
+Specifically, Dirson et al.
+(2022) study was limited by the assumption that pulsars move in a ballistic motion depicted by straight lines at a constant speed from their birth to the present time.
+A noticeable discrepancy arises when comparing the observed position of pulsars to the predictions, revealing a significant disparity in both spatial distributions.
+Dirson et al.
+(2022) did not require a death line to reproduce the $P-\dot{P}$diagram and was therefore not implemented in their PPS.
+However the death line segregates neutron stars generating particles through pair cascading from those not radiating any more because of the quenching of this pair avalanche, and hence they are called dead pulsars.
+The effect of the death line assumption on the PPS differs from the PPS performed by Graber et al.
+(2024) who did not consider a death line but instead relied only on as special radio luminosity law to eliminate some pulsars in the $P-\dot{P}$ diagram.
+In this paper we further improve the PPS of Dirson et al.
+(2022) by examining the necessity of a death line to explain the observed canonical pulsars population and discuss in depth the $\gamma$-ray pulsar population.
+Moreover, our study helps to better understand the impact of changing the parameters of our model as well as their physically meaningful ranges.
 The paper is organised as follows.
 In Sect.
-2 we detail our model, recalling the processes for generation and evolution of the pulsars in the PPS, a description of the Galactic potential, the death line and the multi-wavelength detection criteria.
-Then the results are shown in Sect.
+2 we detail our PPS model, recalling the processes for generation and evolution of the pulsar sample, the description of the Galactic potential and of the death line, discussing our adopted multi-wavelength detection criteria.
+The results are shown in Sect.
 3.
-The paper ends with a discussion in Sect.
-4 and a summary in Sect.
-5.
+The results are discussed in Sect.
+4 and Sect.
+5 covers a summary of the results.
 2 Description of the PPS model
-Our PPS study is based on a self-consistent state-of-the-art evolution of the pulsar geometry and proper motion within the Milky Way.
-In this section we detail the model by exposing first the generation of individual pulsars, their initial position, magnetic field and spin period, then how these quantities should evolve in time, giving explicit expressions for the Galactic potential used in our simulations, before introducing the death line and finally the radio and $\gamma$-ray detection criteria.
+Our PPS study is based on a self-consistent state-of-the-art evolution of the neutron star geometry of radio and $\gamma$-ray emission, magnetic field and proper motion within the Milky Way.
+In this section we detail our model by exposing, first the generation of individual pulsars, their initial position, magnetic field and spin period, second how these quantities evolve in time.
+We also give explicit expressions for the Galactic potential used in our simulations and introduce the death line before discussing the radio and $\gamma$-ray detection criteria.
+We stress that the model parameters used in this section have been carefully chosen by comparing simulated data with actual observations, varying the unknown input parameters in sensible physical ranges (see appendix C for more details).
 2.1 Generation of pulsars
-We generate a population of pulsars whose ages are generated in ascending order, every X years for example if the birth rate is 1/X yr-1 between 0 yr and the age of the Milky Way at maximum.
-The age of the Milky Way is almost equal to the age of the Universe, which is of about $t_{H}\approx 13.8\times 10^{9}$ yr, therefore, a total number of $t_{H}/X$ pulsars should be simulated.
-However as we will see, we chose to emulate $10^{7}$ pulsars in a first run and $10^{6}$ in another run.
-This choice was made for two reasons : firstly, the computation time is significantly decreased while pulsars older than $10^{8}$ yr are mostly not detected, secondly, the simulation that emulates $10^{7}$ pulsars allows to emulate older pulsars and require a death line while the other works well without.
-This approach was also chosen because it allows to check whether our birth rate is relevant by comparing our results from the simulation to the observations.
-It is different from other works like Johnston & Karastergiou (2017) and Gullon et al.
-(2014), where they stop the simulation when the number of pulsars detected in the simulation are the same as the number of observed pulsars.
-The value chosen for the birth rate, which is 1/41 yr-1, is taken as best educated guess from the interval [1/150,1/33] yr-1 because these studies (Faucher-Giguère & Kaspi, 2006; Gullon et al., 2014; Johnston & Karastergiou, 2017) evaluated the birth rate of pulsars in this interval.
+We generate a population of pulsars whose real ages are selected in ascending order, for example every X years if the birth rate is 1/X yr-1, with their age being between 0 yr and the age of the Milky Way at maximum.
+The age of the Milky Way does not significantly differ from the age of the Universe estimated about $t_{H}\approx 13.8\times 10^{9}$ yr, therefore, a total number of $t_{H}/X$ pulsars should be simulated.
+The best value for this birth rate was found to be 1/41 yr-1, see appendix C.
+Our value is consistent with earlier investigations by Faucher-Giguère & Kaspi (2006); Gullon et al.
+(2014); Johnston & Karastergiou (2017) where they evaluated the birth rate of pulsars to be in the range [1/150,1/33] yr-1.
+With $X=41$ yr we should generate more than $10^{7}$ pulsars, however we chose to emulate only $10^{7}$ pulsars in a first run and $10^{6}$ in a second run for the following two reasons.
+Firstly, the computation time is significantly decreased, secondly pulsars older than $10^{7}$ yr are mostly not detected.
+This approach was also chosen because it allows us to verify whether our birth rate is realistic, by comparing our results from the simulation to the observations.
+Our approach is different from previous works like Gullon et al.
+(2014) and Johnston & Karastergiou (2017), where they stop the simulation whenever their detected number of pulsars equals the number of observed pulsars.
 To describe the position of the pulsars in the Galaxy, we use the right-handed Galacto-centric coordinate system $(x,y,z)$ with the Galactic centre at its origin, $y$ increasing in the disc plane towards the location of the Sun, and $z$ increasing towards the direction of the north Galactic pole.
 The initial spatial distribution of the pulsars is given by Paczynski (1990).
-He found two distributions, one describing the radial spread and one the spread in altitude, however only his distribution in altitude is used in our work because the radial distribution in the galactic plane is better depict in Ahlers et al.
-(2016) by the Milky Way’s pulsar surface density defined as
+He found two distributions, one describing the radial spread and one the spread in altitude, however only his distribution in altitude is used in our work because the radial distribution in the galactic plane is better depicted in Yusifov & Küçük (2004) by the Milky Way’s pulsar surface density defined as
 $\rho(R)$ $=A\left(\frac{R+R_{1}}{R_{\odot}+R_{1}}\right)^{a}\exp\left(-b\left(\frac{R-R_{\odot}}{R_{\odot}+R_{1}}\right)\right)$ (1a)
@@
 where $R$ is the axial distance from the $z$-axis, and $z$ is the distance from the Galactic disc.
-The numerical values for the constants are, $A=37.6$ kpc-2, $a=1.93$, $b=5.06$, $R_{1}=0.55$ kpc, $h_{c}=180$ pc and $R_{\odot}=8.5$ kpc and in agreement with the distribution of young massive stars in our galaxy (Li et al., 2019).
+The numerical values for the constants are, $A=37.6$ kpc-2, $a=1.64$, $b=4.0$, $R_{1}=0.55$ kpc, $h_{c}=180$ pc and $R_{\odot}=8.5$ kpc and in agreement with the distribution of young massive stars in our galaxy (Li et al., 2019).
 The purpose of choosing equation (1a) to describe the radial spread is to put neutron stars birth positions within the spiral arms of the Galaxy.
 More precisely the Galactic spiral structure contains four arms with a logarithmic shape function allowing to obtain the azimuthal coordinate $\phi$ as a function of the distance from the Galactic center
-$\phi(R)=k\ln\left(\frac{R}{r_{0}}\right)+\phi_{0}$ (2)
+$\phi(R)=k\ln\left(\frac{R}{r_{0}}\right)+\phi_{0}.$ (2)
 The values of the models describing each arm are given in Table 2.
@@
 As soon as $R_{birth}$ and $\phi_{birth}$ are known we can convert these coordinates in $x$ and $y$ Galacto-centric coordinates for each pulsar.
-Table 2: Parameters of the Milky Way Spiral Arm structure.
+Table 2: Parameters of the Milky Way Spiral Arm structure, adapted from Yao et al. (2017).
 Arm Number Name $k$ $r_{0}$ $\phi_{0}$
@@
 4 Crux-Scutum 5.37 3.67 5.76
-Next, we need to describe the individual properties of each pulsar, namely, its inclination angle $\alpha$, which is the angle between its rotation axis and its magnetic axis, its initial spin period and initial magnetic field at birth.
+Next, we need to describe the individual properties of each pulsar, namely, its inclination angle $\alpha$, which is the angle between its rotation axis and its magnetic axis, its initial spin period $P_{0}$ and initial magnetic field $B_{0}$ at birth.
 The inclination angle $\alpha$ is assumed to follow an isotropic distribution generated from a uniform distribution $U\in[0,1]$ and given by $\alpha=\arccos(2\,U-1)$.
 In our work, the initial spin period and magnetic field follow both a log-normal distribution, as suggested from the results of a study of 56 young neutron stars by Igoshev et al.
-(2022) while other works simply use a Gaussian distribution (Faucher-Giguère & Kaspi, 2006; Gullon et al., 2014; Johnston et al., 2020).
-Explicitly, the probabilities to born with an initial magnetic field $B_{0}$ and period $P_{0}$ are given by
+(2022) while other works simply use a Gaussian distribution for the spin period at birth (Faucher-Giguère & Kaspi, 2006; Gullon et al., 2014; Johnston et al., 2020) while for the magnetic field the prescription is the same as ours.
+Explicitly, the probability distributions for an initial magnetic field $B_{0}$ and an initial period $P_{0}$ are given by
 $p(\log(B_{0}))=\frac{1}{\sigma_{b}\sqrt{2\pi}}e^{-(\log B_{0}-\log\bar{B})^{2}/(2\sigma_{b}^{2})}$ (3)
@@
 (2020), Yadigaroglu & Romani (1995) and Watters & Romani (2011).
-Compared to the study of Dirson et al.
+According to Hobbs et al. (2005) a Maxwellian distribution for the kick velocity at birth best replicates the observations and is given by
+$p(v)=\sqrt{\frac{2}{\pi}}\frac{v^{2}}{\sigma_{v}^{3}}\exp{\left(-\frac{v^{2}}{2\sigma_{v}^{2}}\right)}$ (5)
+The mean velocity of the distribution is related to the standard deviation by $\bar{v}=\sigma_{\rm v}\sqrt{{8}/{\pi}}$ with a standard deviation $\sigma_{\rm v}=265$ km/s for pulsars.
+The velocity value drawn from the Maxwellian distribution is then distributed along the unit vector of the rotation axis which is generated, as detailed in subsection 2.5.
+Moreover compared to the study of Dirson et al.
 (2022) we introduce a novelty, namely the alignment between the kick velocities at birth and the rotation axis of the pulsar, as suggested by Rankin (2007).
-Moreover according to Hobbs et al.
-(2005) a Maxwellian distribution best replicates the observations, with a standard deviation $\sigma_{\rm v}=265$ km/s.
-The mean velocity of the distribution is related to the standard deviation by $\bar{v}=\sigma_{\rm v}\sqrt{{8}/{\pi}}$.
-The velocity vector is then distributed along the unit vector of the rotation axis generated as detailed in the subsection 2.5.
 2.2 Pulsar evolution
-The pulsar initial period, inclination angle and magnetic field must be evolved in time self-consistently due to electromagnetic radiation and internal dissipation within the neutron star.
-To this end we employ the pulsar force-free magnetosphere model taking into account the plasma current and charge within the magnetosphere.
+The pulsar initial period $P_{0}$, inclination angle $\alpha_{0}$ and magnetic field $B_{0}$ are evolved in time in a fully self-consistent way taking into account the spin-down losses and the internal magnetic field dissipation within the neutron star.
+The neutron star period is evolved according to the force-free magnetosphere model which takes into account the electric current and charge flowing inside the magnetosphere.
 The spin down luminosity $\dot{E}$ is
@@
 $L_{\rm ffe}=\frac{4\pi R^{6}B^{2}\Omega^{4}(1+\sin^{2}\alpha)}{\mu_{0}c^{3}}$ (7)
-where $I$ is the neutron star moment of inertia its value being approximately $I\approx 10^{38}$ kg.m2, $\Omega=2\pi/P$ is the rotation frequency of the pulsar ($P$ is the spin period) and $\dot{\Omega}$ is its time derivative, $R=12$ km is the typical radius of a neutron star as found by recent NICER observations (Riley et al., 2019; Bogdanov et al., 2019), $B$ is the magnetic field at the magnetic equator, $\alpha$ is the inclination angle, $\mu_{0}$ is the vacuum permeability constant, $\mu_{0}=4\pi\times 10^{-7}$ H/m and $c$ is the speed of light.
+where $I$ is the neutron star moment of inertia approximately equal to $I\approx 10^{38}$ kg.m2, $\Omega=2\pi/P$ the rotation frequency of the pulsar ($P$ is the spin period), $\dot{\Omega}$ its time derivative, $R=12$ km the typical radius of a neutron star as found by recent NICER observations (Riley et al., 2019; Bogdanov et al., 2019), $B$ the magnetic field at the magnetic equator, $\alpha$ the inclination angle, $\mu_{0}$ the vacuum permeability constant, $\mu_{0}=4\pi\times 10^{-7}$ H/m and $c$ the speed of light.
 The term $L_{\rm ffe}$ for the spin down luminosity was given by Spitkovsky (2006) and Pétri (2012).
@@
 $\Omega\frac{\cos^{2}\alpha}{\sin\alpha}=\Omega_{0}\frac{\cos^{2}\alpha_{0}}{\sin\alpha_{0}}$ (10)
-where quantities with index 0 indicate their initial value. The quantities without the index are their current time value.
+where quantities with subscript 0 indicate their initial value and those without a subscript their current value at present time.
 As shown by Gullon et al.
@@
 $B(t)=B_{0}(1+t/\tau_{d})^{-1/\alpha_{d}}$ (11)
-where $\alpha_{d}$ is a constant parameter controlling the speed of the magnetic field decay and $\tau_{d}=k_{\tau_{d}}\tau_{v}$ is the magnetic field decay time scale.
-The decay rate $\tau_{v}$ was chosen randomly between 3 values : $5\times 10^{4}$ yr or $7\times 10^{4}$ yr or $5\times 10^{5}$ yr, with a probability of 0.23, 0.46 and 0.31 respectively (these probabilities were chosen as best educated guesses after several run with different values) for each pulsar, corresponding to the $\tau_{v}$ value for canonical pulsars in Viganò et al.
-(2013).
-The values of $\alpha_{d}$ and $k_{\tau_{d}}$ can be found in Table 5 of Sect.
+where $\alpha_{d}$ is a constant parameter controlling the rate of the magnetic field decay and $\tau_{d}$ the typical decay timescale, which depends on the initial magnetic field as demonstrated by magneto-thermal evolution models (Viganò et al., 2013).
+We took these results into account by defining the decay timescale $\tau_{1}$ for a magnetic field strength $B_{1}$ such that $\tau_{1}\,B_{1}^{\alpha_{d}}=\tau_{d}\,B_{0}^{\alpha_{d}}$.
+Thus, $\tau_{1}$ is another magnetic field decay time scale, corresponding to a certain magnetic field value $B_{1}$.
+Three different $\tau_{1}$ are considered in this work and are randomly chosen for each pulsar generated : $1.5\times 10^{5}$ yr or $3.5\times 10^{5}$ yr or $2.5\times 10^{6}$ yr with each of these timescales corresponds to a magnetic field value $B_{1}$ of $1\times 10^{8}$ T, $3\times 10^{8}$ T and $2\times 10^{9}$ T (these values of $\tau_{1}$ and $B_{1}$ are taken from Fig.
+10 of the work of Viganò et al.
+(2013)).
+Then the probabilities to get these values of $\tau_{1}$ and $B_{1}$ to be adopted for one pulsar are 0.23, 0.46 and 0.31 respectively (these probabilities were chosen as best educated guesses after several run with different values).
+The value of $\alpha_{d}$ can be found in Table 5 of Sect.
 3.
@@
 (2014) for a spherically symmetric neutron star with a constant magnetic field.
-With our decaying prescription the inclination angle $\alpha$ is found by solving for the root of
+With our decaying prescription the inclination angle $\alpha$ is found by solving
 $\begin{split}\ln(\sin\alpha_{0})+\frac{1}{2\sin^{2}\alpha_{0}}+K\Omega_{0}^{2}\frac{\cos^{4}\alpha_{0}}{\sin^{2}\alpha_{0}}\frac{\alpha_{d}\tau_{d}B_{0}^{2}}{\alpha_{d}-2}\left[\left(1+\frac{t}{\tau_{d}}\right)^{1-2/\alpha_{d}}-1\right]\\ \phantom{{}={}}=\ln(\sin\alpha)+\frac{1}{2\sin^{2}\alpha}\end{split}$ (12)
 where $t$ is the time representing the age of the pulsar and $K={R^{6}}/{I\,c^{3}}$.
-The typical decay timescale for a mean magnetic field of $2.5\times 10^{8}$ T is $4.6\times 10^{5}$ yr in Viganò (2013), however in this study the typical decay timescale associated to this magnetic field is $3.5\times 10^{5}$ yr which is consistent with the estimate of Viganò (2013) for the same field strength.
-Finally, we need to follow the particle motion within the Galactic potential. Each generated pulsar evolves in the gravitational potential $\Phi$ subject to an acceleration ${\ddot{x}}$ according to
+The typical decay timescale for a mean magnetic field of $2.5\times 10^{8}$ T is $4.6\times 10^{5}$ yr in Viganò (2013), however in this study the typical decay timescale associated to this magnetic field is $3.5\times 10^{5}$ yr which is very close to the estimate of Viganò (2013) for the same field strength.
+2.3 Description of the Galactic potential
+Finally, we need to follow the particle motion within the Galactic potential. Each pulsar evolves in the gravitational potential $\Phi$ subject to an acceleration ${\ddot{x}}$ according to
 ${\ddot{x}}=-\@vec{\nabla}\Phi\ .$ (13)
-Equation (12) is integrated numerically thanks to a Position Extended Forest Ruth-Like (PEFRL) algorithm (Peter Young, 2014), a fourth order integration scheme, presented in Appendix A with a convergence and accuracy study.
-The galactic potential model used in discussed in the next section.
-2.3 Description of the Galactic potential
+Equation (13) is integrated numerically thanks to a Position Extended Forest Ruth-Like (PEFRL) algorithm (Young, 2014), a fourth order integration scheme, presented in Appendix A with a convergence and accuracy study.
+We now discuss the galactic potential model used.
 The Galaxy is divided in four distinct regions with different mass distributions and associated gravitational potentials.
@@
 Equation (17) is the derivative of the Miyamoto Nagai potential with respect to the $z$ coordinate.
-The potential for the dark matter halo $\Phi_{h}$, according to the frequently used potential of Navarro et al. (1997), it is expressed as
+The potential for the dark matter halo $\Phi_{h}$, according to the frequently used potential of Navarro et al. (1997), is expressed as
 $\Phi_{h}(r)=-\frac{GM_{h}}{r}\ \ln\left(1+\frac{r}{a_{h}}\right)$ (18)
@@
 The most suitable one in our study is that of Mitra et al.
-(2019) where they found an expression for the death line
+(2019) where they found an expression relating $P$ and $\dot{P}$ by
 $\dot{P}_{\rm line}=\frac{3.16\times 10^{-19}\ T_{6}^{4}\ P^{2}}{\eta^{2}b\cos^{2}\alpha_{l}}\ .$ (22)
@@
 Spread in the parameter values of the model causes significant variations in the death line, and thus a death valley rather than a single death line describes the condition for pulsar extinction in the $P-\dot{P}$ diagram.
-To decide whether a pulsar is still active in radio, its critical spin down rate $\dot{P}_{\rm line}$ is computed knowing its spin period $P$.
-If the pulsar lies in the death valley defined by the inequality $10^{-0.55}<\dot{P}_{\rm line}/\dot{P}<10^{1.15}$ then a value for $\alpha_{l}$ is taken from a uniform distribution between $0\degr$ and $65\degr$.
-Moreover a value for $T_{6}$ is also taken from a uniform distribution between 1.9 and 2.8.
-The parameter $b$ is taken from a uniform distribution between 30 and 60.
-The boundary parameter values for the death valley were chosen to traverse the $P-\dot{P}$ diagram from the region in the bottom right quadrant where the density starts to decrease, to the detectable pulsar with the lowest $\dot{P}$ and highest $P$ in the same quadrant, as shown in Fig.
+We allow the different parameters for this death valley to be in the following ranges : $\alpha_{l}$ is drawn from a uniform distribution between $0\degr$ and $65\degr$, $T_{6}$ is drawn from a uniform distribution between 1.9 and 2.8 and $b$ is drawn from a uniform distribution between 30 and 60.
+The boundary parameter values, corresponding to the edges of the death valley, were chosen to traverse the $P-\dot{P}$ diagram from the region in the bottom right quadrant where the number of pulsars starts to decrease (because this is likely the region where the death of pulsars begins to be considered, since less pulsars are there), to the detectable pulsar with the lowest $\dot{P}$ and highest $P$ in the same quadrant, as shown in Fig.
 1.
-The edges of the death valley given by the values $10^{-0.55}$ and $10^{1.15}$ were found by taking the extreme values for the parameters that are drawn from the different uniform distributions.
-Then we compare its actual spin down rate $\dot{P}$, computed from the evolution model, to the $\dot{P}^{\rm new}_{\rm line}$ computed with equation (21) with the parameters which were drawn from uniform distributions.
+To decide whether a pulsar is still active in radio or not we apply the following criterion: if the pulsar lies in the death valley then all the parameters mentioned above are taken from uniform distributions in the ranges of the death valley in order to compute $\dot{P}^{\rm new}_{\rm line}$ with equation (22), the critical spin down rate of the considered pulsar.
+Then we compare its actual spin down rate $\dot{P}$, computed from the evolution model, to the $\dot{P}^{\rm new}_{\rm line}$.
 For pulsars that are not in the death valley, their $\dot{P}$ is only compared with the spin down rate $\dot{P}_{\rm line}$ obtained with the default parameters.
-If the pulsar has its $\dot{P}$, which was computed from the evolution model, below the one computed with the death line ($\dot{P}_{\rm line}$ or $\dot{P}^{\rm new}_{\rm line}$), then the pulsar is considered as dead, otherwise it is considered as still emitting photons.
+If the pulsar has its $\dot{P}$, which was computed from the evolution model, below the one computed with the death line, then the pulsar is considered as dead, otherwise it is considered as still emitting photons.
 These choices were made in order to have more like a death valley near the death line than a strict condition of death, because the parameters $T_{6}$, $\alpha_{l}$ and $b$ may be different from one pulsar to another.
@@
 2.5 Detection
-For each generated pulsar, we check whether it fulfills the criteria for detection depending on three factors.
+For each pulsar, we check whether it fulfills the detection criteria depending on three factors.
 First, the beaming fraction indicates the fraction of the sky covered by the radiation beam and depends on the considered wavelength, here radio or $\gamma$-ray.
@@
 Moreover, it is related to the previous angles by $\alpha+\beta=\xi$.
-We choose a isotropic distribution for the Earth viewing angle $\xi$ as well as for the orientation of the unitary rotation vector.
+We choose an isotropic distribution for the Earth viewing angle $\xi$ as well as for the orientation of the unitary rotation vector.
 The Cartesian coordinates of the unit rotation vector ${n_{\Omega}}$ are $(\sin\theta_{n_{\Omega}}\cos\phi_{n_{\Omega}},\sin\theta_{n_{\Omega}}\sin\phi_{n_{\Omega}},\cos\theta_{n_{\Omega}})$.
@@
 2.5.1 Radio detection model
-The beaming fraction in radio depends on the half opening angle of the radio emission cone $\rho$ computed according to
+The beaming fraction in radio depends on the half opening angle of the radio emission cone $\rho$ computed according to Lorimer & Kramer (2004) by
 $\rho=3\sqrt{\frac{\pi\,h_{\rm em}}{2Pc}}$ (25)
 where $h_{\rm em}$ is the emission height, $P$ is the spin period of the pulsar and $c$ is the speed of light.
-The emission height is taken constant with an average value of $h_{\rm em}=3.10^{5}$ m, estimated from observations of a sample of pulsars by (Weltevrede & Johnston, 2008; Mitra, 2017; Johnston & Karastergiou, 2019; Johnston et al., 2023).
+The emission height is taken constant with an average value of $h_{\rm em}=3.10^{5}$ m, estimated from observations of a sample of pulsars by Weltevrede & Johnston (2008); Mitra (2017); Johnston & Karastergiou (2019); Johnston et al.
+(2023).
 The cone half-opening angle $\rho$ estimated in eq.
@@
 $\cos(\rho)=\cos(\alpha)\cos(\xi)+\sin(\alpha)\sin(\xi)\cos(w_{r}/2).$ (26)
-The second factor for detection is the luminosity, which is also different between radio and gamma-rays.
+The second factor for detection is the luminosity, which is also different between radio and $\gamma$-rays.
 First, concerning the radio flux density, the formula used is the same as in Johnston et al.
-(2020) for a pulsar at 1.4 GHz, in order to model the detection carried out by the Parkes radio telescope in the Southern Galactic plane (Kramer et al., 2003; Lorimer et al., 2006; Cameron et al., 2020) and the Arecibo telescope in the northern plane (Cordes et al., 2006)
+(2020) for a pulsar at 1.4 GHz, in order to model the detection carried out by the Parkes radio telescope in the southern Galactic plane (Kramer et al., 2003; Lorimer et al., 2006; Cameron et al., 2020) and the Arecibo telescope in the northern plane (Cordes et al., 2006) (note that these two telescopes do not cover all the sky in the southern and northern plane, but we did not implemented any filter in the code)
 $F_{r}=9\ \text{mJy}\left(\frac{d}{1\ \text{kpc}}\right)^{-2}\left(\frac{\dot{E}}{10^{29}\ \text{W}}\right)^{1/4}\times 10^{F_{\rm j}}$ (27)
@@
 With the aim of computing the observed width pulse profile, we use the same formula as Cordes & McLaughlin (2003)
-$\tilde{w}_{r}=\sqrt{\left(w_{r}\frac{P}{2\pi}\right)^{2}+\tau_{samp}^{2}+\tau_{DM}^{2}+\tau_{scat}^{2}}$ (28a)
+$\tilde{w}_{r}=\sqrt{\left({w_{r}\,P}/{2\pi}\right)^{2}+\tau_{samp}^{2}+\tau_{DM}^{2}+\tau_{scat}^{2}}$ (29a)
 $S_{\text{survey}}^{\text{min}}=S_{0}\sqrt{\frac{\tilde{w}_{r}}{P-\tilde{w}_{r}}}$ (29b)
@@
 (2017) converting the distance of a pulsar into the dispersion measure thanks to their state-of-the art model of the Galactic electron density distribution.
-The radio survey parameters used are the ones from the Parkes Multibeam Pulsar Survey (PMPS), see Table 4.
+The radio survey parameters used are the ones from the Parkes Multibeam Pulsar Survey (PMPS) taken from Manchester et al.
+(2001), see Table 4.
 Table 4: Survey parameters of the Parks Multibeam Pulsar Survey (PMPS).
@@
 In addition we only use the parameters of PMPS to compute $\tau_{DM}$, we could have also used the parameters of The Pulsar Arecibo L-band Feed Array (PALFA) Survey, however the differences in the results in the end are not large if we use these parameters, consequently, it persists as a favorable approximation to exclusively utilize the parameters delineated by PMPS.
-2.5.2 Gamma-ray detection model
-The $\gamma$-ray emission model relies on the striped wind model, describing gamma-ray photons production emanating from the current sheet within the striped wind.
-In order to detect the gamma-ray, the line of sight of the observer must remain around the equator plane with an inclination angle constrained by $|\xi-\pi/2|\leq\alpha$.
-The gamma-ray luminosity is extracted from a study of Kalapotharakos et al.
+2.5.2 $\gamma$-ray detection model
+The $\gamma$-ray emission model relies on the striped wind model, describing $\gamma$-ray photons production emanating from the current sheet within the striped wind.
+In order to detect the $\gamma$-ray, the line of sight of the observer must remain around the equator plane with an inclination angle constrained by $|\xi-\pi/2|\leq\alpha$.
+The $\gamma$-ray luminosity is extracted from a study of Kalapotharakos et al.
 (2019), where they showed that the luminosity is described by a fundamental plane.
@@
 For the striped wind model Pétri (2011) showed that this factor varies between 0.22 and 1.90.
-Nevertheless, an approximation is made : if $\alpha<-\xi+0.6109$ then $f_{\Omega}=1.9$ otherwise $f_{\Omega}=1$.
+Nevertheless, an approximation is made : if $\alpha<-\xi+0.6109$ then $f_{\Omega}=1.9$ otherwise $f_{\Omega}=1$, since as can be observed in Fig.
+7 of Pétri (2011) $f_{\Omega}$ is usually equal to one of this two values depending on the value of the inclination angle $\alpha$, hence this approximation.
 The pulsar is detected in gamma depending on the instrumental sensitivity as described below.
-The sensitivity in gamma-ray is based on the expectation of the Fermi/LAT instrument  [^4] https://fermi.gsfc.nasa.gov/ssc/data/analysis/documentation/Cicerone
+The sensitivity in $\gamma$-ray is based on the expectation of the Fermi/LAT instrument  [^5] https://fermi.gsfc.nasa.gov/ssc/data/analysis/documentation/Cicerone
 /Cicerone_LAT_IRFs/LAT_sensitivity.html
@@
 Let us now move to the PPS simulation results.
-First we show runs simulating 10 million pulsars with and without a death line, allowing to simulate old pulsars (with the oldest aged $4.1\times 10^{8}$ yr).
-Then we show runs simulating 1 million pulsars (with the oldest aged $4.1\times 10^{7}$ yr) with and without considering a death line.
-We explore a possible spin-velocity misalignment effect due to the galactic potential and the impact of the initial spin period distribution, normal vs log-normal.
-Finally, the gamma-ray simulations and detection are discussed in details.
+First we show runs simulating 10 million pulsars with and without a death line, allowing to generate old pulsars with ages up to $4.1\times 10^{8}$ yr.
+Then we show runs simulating only 1 million pulsars with ages up to $4.1\times 10^{7}$ yr, also with and without a death line.
+We explore a possible spin-velocity misalignment effect due to the galactic potential and the impact of the initial spin period distribution, comparing normal vs log-normal distributions.
+Finally, the $\gamma$-ray simulations and detection are exposed in details.
 3.1 Simulating 10 millions pulsars
+Let us start with a sample containing 10 millions pulsars showing the impact of the death line. A same study will be done for 1 million pulsars.
 3.1.1 Without a death line
-The first simulations performed with the Galactic potential implemented discarded the death line.
-The parameters used for these runs are given in Table 5.
-The total number of pulsars is summarized in Table 6.
+The first set of simulations discarded the death line. The parameters used for these runs are given in Table 5. The total number of pulsars is summarized in Table 6.
 Table 5: Parameters used in the simulations.
-$\tau_{\text{birth}}$ (1/yr) $P_{\text{mean}}$ (ms) $B_{\text{mean}}$ (T) $\sigma_{p}$ $\sigma_{b}$ $\alpha_{d}$ $k_{\tau_{d}}$
-41 129 $2.75\times 10^{8}$ 0.45 0.5 1.5 5
+$\tau_{\text{birth}}$ (1/yr) $P_{\text{mean}}$ (ms) $B_{\text{mean}}$ (T) $\sigma_{p}$ $\sigma_{b}$ $\alpha_{d}$
+41 129 $2.75\times 10^{8}$ 0.45 0.5 1.5
 Table 6: Number of pulsars detected without the implementation of the death line.
 log($\dot{E}$) (W) $N_{tot}$ $N_{r}$ $N_{g}$ $N_{rg}$
-¿31 11 0 8 3
-¿28 161 9 75 77
-Total 2653 2336 139 178
-The quantities $N_{r}$, $N_{g}$, and $N_{rg}$ are the number of radio-only, gamma-only, and radio-loud gamma-ray pulsars, respectively, obtained from our simulation.
+¿31 7 0 4 3
+¿28 161 4 87 63
+Total 2491 2169 163 159
+The quantities $N_{r}$, $N_{g}$, and $N_{rg}$ are the number of radio-only, $\gamma$-only, and radio-loud $\gamma$-ray pulsars, respectively, extracted from our simulation.
 Compared to the results obtained by Dirson et al.
-(2022), at very high spin down luminosity, $\dot{E}>10^{31}$ W, the detection is almost identical but at $\dot{E}>10^{28}$ W there are much less pulsars detected in radio and radio/gamma in this work.
-There are probably two causes for this discrepancy in the number of detected pulsars at $\dot{E}>10^{28}$ W : firstly the difference in the spin period distribution at birth.
-Pulsars have a higher probability to start with a higher $P_{0}$ since the mean of the distribution is higher (in the previous study the mean was 60 ms while it is 129 ms now), therefore the slowing down of the pulsars can only decrease their rotation frequency $\Omega$, and equation (5) shows that $\dot{E}$ depends on $\Omega$.
-However, it does not affect the detection in gamma-rays because usually canonical gamma pulsars have a high $\dot{P}$ and a low $P$, being young, therefore their $\dot{E}$ is higher than for radio pulsars.
-Secondly, the ISM dispersion and scattering reduces this number of radio pulsars detected (the results without the ISM effect are not shown but without it there is much more radio pulsars detected).
-Nonetheless, in terms of total number of detected radio pulsars, the Galactic potential allows to have an increase compared to the number found in Dirson et al.
+(2022), at very high spin down luminosity, $\dot{E}>10^{31}$ W, the detection rate is almost identical but at $\dot{E}>10^{28}$ W much less pulsars are detected in radio and radio/$\gamma$-ray in our work.
+There are probably two causes for this discrepancy in the number of detected pulsars at high spin down luminosities $\dot{E}>10^{28}$ W : firstly the difference in the spin period distribution at birth.
+Pulsars have a higher probability to start with a higher $P_{0}$ since the mean of the distribution is higher (in the previous study the mean period was 60 ms while we set it to 129 ms), therefore the slowing down of the pulsars can only decrease their rotation frequency $\Omega$, and equation (6) shows that $\dot{E}$ depends on $\Omega$.
+However, it does not affect the detection rate in $\gamma$-rays because usually canonical $\gamma$-ray pulsars possess a high $\dot{P}$ and a low $P$, being young, therefore their $\dot{E}$ is higher than for radio pulsars.
+Secondly, the ISM dispersion and scattering reduces the number of detected radio pulsars, results without the ISM effect are not shown but discarding its impact would lead to much more detected radio pulsars.
+Nonetheless, in terms of total number of detected radio pulsars, including the Galactic potential increases this number compared to Dirson et al.
 (2022), a result expected because of the attractive nature of the potential of the Galaxy.
-In its previous version the PPS moved the pulsars in random spatial directions at constant speed, as a result no pulsar was bound to the Milky Way in closed orbits this is especially true for old pulsars which could have time to leave the Galaxy.
+Moreover in this latter version, the PPS moved pulsars in random spatial directions at constant speed, as a result no pulsar was bound to the Milky Way in closed orbits an artifact that is especially true for old pulsars which could have time to leave the Galaxy.
 Figure 2: Distribution of the observed period taken from the ATNF catalogue, along with the simulations without the implementation of the death line.
-In green, the period of the simulated pulsars with age greater than $10^{7.5}$
+In green, the period of the simulated pulsars with characteristic age greater than $10^{8}$ yr.
 Fig.
-2 shows the distribution of the spin period of the detected pulsars in the simulation, in red and in observations, in blue.
-However the comparison is not satisfactory because there is an excess of pulsars detected in the simulation.
+2 compares the distribution of simulated and observed spin periods, in red and blue respectively.
+However the comparison is not satisfactory because of an excess of pulsars detected in the simulation.
 Furthermore, this excess is also present in Fig.
-3 where the age of the detected pulsars from the observations and simulation are plotted.
-In order to verify if there is a correlation between old pulsars and the ones with a high spin period, we plotted in Fig.
-2, in green the pulsars with an age greater than $10^{7.5}$ yr, corresponding to the oldest pulsars of the simulation and it corresponds to the excess found in Fig.
+3 where the characteristic age of the observed pulsars and the simulated pulsars are plotted.
+Note that we employ the characteristic ages defined by
+$\tau_{c}=\frac{P}{2\dot{P}}$ (34)
+because we usually do not know the real ages of the observed pulsars.
+Figure 3: Distribution of the observed age taken from the ATNF catalogue, along with the simulations without the implementation of the death line.
+In order to verify whether there is a correlation between old pulsars and those with a high spin period or not, Fig.
+2 shows in green pulsars with a characteristic age greater than $10^{8}$ yr, corresponding to the oldest pulsars of the simulation.
+It coincides with the excess found in Fig.
 2.
-Due to the Galactic potential, more pulsars remain bound to the Galaxy, especially older pulsars which were escaping the Galaxy more easily in the work of Dirson et al.
-(2022) because of the absence of potential.
+We conclude that canonical pulsars older than $10^{8}$ yr measured as their real age in our simulations should be discarded at the end of the simulation, since the characteristic age is overestimating the real age of pulsars.
+Indeed canonical pulsars older than this do not exist.
+The Galactic potential forbids many pulsars to escape from the Milky Way, they remain bound to the Galaxy, in contrast to the older pulsars found in Dirson et al.
+(2022), that were escaping the Galaxy easily because of the absence of this potential.
 Fig.
-4 shows that many pulsars are in what is commonly called the graveyard of pulsars at the bottom right of the $P-\dot{P}$ diagram.
-Figure 3: Distribution of the observed age taken from the ATNF catalogue, along with the simulations without the implementation of the death line.
+4 shows that many pulsars are in what is commonly called the graveyard of pulsars at the bottom right of the $P-\dot{P}$ diagram, the death valley being shown in Fig.
+4 but not implemented.
 Figure 4: $P-\dot{P}$ diagram of the simulated population, along with the observations, without the implementation of the death line.
-Given this excess of long period pulsars without a death line, we decided to implement one, in order to stick closer to reality.
+Given this excess of long period pulsars when neglecting the death line, we decided to implement it, in order to stick closer to reality.
 3.1.2 With the death line
+Figure 5: Distribution of the observed period taken from the ATNF catalogue, along with the simulations with the death line implementation.
 Adding a death line decreases the number of pulsars detected in radio compared to the situation without a death line, see Table 7.
+The death line improves the fit as seen in Fig.
+5 for the distribution of spin period and compared with Fig.
+2, even though there is a slight excess of pulsars which have spin period between $10^{-0.5}$ s and $1$ s.
+In Fig.
+6, these pulsars in excess lie at $\dot{P}$ between $10^{-18}$ and $10^{-17}$, where there is no data and which correspond to the oldest pulsars in the simulation.
+Therefore the oldest pulsars simulated here do not correspond to any real data.
+In addition, too many pulsars end up close to the death line, disagreeing with the observations.
+The death line creates a pile up effect not seen in the observations.
 Table 7: Number of pulsars detected with the death line implementation.
@@
 Total 2077 1807 116 154
-Figure 5: Distribution of the observed period taken from the ATNF catalogue, along with the simulations with the death line implementation.
 Figure 6: $P-\dot{P}$ diagram of the simulated population, along with the observations, with the death line implementation.
-Figure 7: Density plot of the $P-\dot{P}$ diagram in comparison with observations
-The death line improves the fit as seen in Fig.
-5 for the distribution of spin period and compared with Fig.
-2, even though there is a slight excess of pulsars which have spin period between $10^{-0.5}$ s and $1$ s.
-In Fig 6, these pulsars in excess lie at $\dot{P}$ between $10^{-18}$ and $10^{-17}$, where there is no data and which correspond to the oldest pulsars in the simulation.
-Therefore the oldest pulsars simulated here do not correspond to any real data.
-In addition, there are too many pulsars close to the death line compared to the observations, there is a pile up effect created by the death line which is not in the observations.
-The goodness-of-fit can be checked using a density plot of the $P-\dot{P}$ diagram as in Johnston & Karastergiou (2017) instead of putting points for individual pulsars.
+Figure 7: Density plot of the $P-\dot{P}$ diagram in comparison with observations.
+The goodness-of-fit is checked against a density plot of the $P-\dot{P}$ diagram as in Johnston & Karastergiou (2017) instead of putting points for individual pulsars.
 In order to construct such diagrams we binned the data evenly in a log scale in $P$ and $\dot{P}$ and count the number of pulsars in each bin.
@@
 $N_{sim}$ is the number of pulsar count in a bin of the simulation, $N_{obs}$ the number of pulsar detected in a bin of the observations.
-A good fit in a bin correspond to a value of $R$ close to 00, meaning that the simulation results are close to the observations.
+A good fit in a bin corresponds to a value of $R$ close to 00, meaning that the simulation results are close to the observations.
 Inspecting fig.
-7, we note a stronger deviation from the observations at the boundaries of the density plot, for instance it confirms too many pulsars are piling up close to the death line.
-The high values of $R$ at these boundaries originates from the weak number of pulsars counted in this bins leading to bad statistics and large variations in the proportions, a side effect of the density plot approach.
+7, we note a stronger deviation from the observations at the boundaries of the density plot, confirming for instance the pile up close to the death line.
+The high values of $R$ at these boundaries originates from the weak number of pulsars counted in these bins leading to bad statistics and large variations in the proportions, a side effect of the density plot approach.
 The fact that the colors are very dark, shows that the proportion of pulsars in most of the bins are not good.
-A Kolmogorov Smirnov (KS) test (Smirnov, 1948) was conducted on both $P$ and $\dot{P}$ which gave p-values substantially below 0.05 respectively when no death line is implemented.
-While we obtained p-values of 0.05 and another one well below 0.05 respectively for $P$ and $\dot{P}$ with a death line implementation.
-The null hypothesis that observations and simulations originate from the same distribution for the spin period can not be rejected (p-value $\geq$ 0.05) only when a death line is considered, but is clearly rejected for $\dot{P}$ in both case.
+A Kolmogorov Smirnov (KS) test (Smirnov, 1948) was conducted on both $P$ and $\dot{P}$ which gave p-values substantially below 0.05 for both quantities, when no death line is implemented.
+While we obtained p-values of 0.05 for $P$ and well below 0.05 for $\dot{P}$ with a death line implementation.
+The null hypothesis that observations and simulations originate from the same distribution for $P$ can not be rejected (p-value $\geq 0.05$) only when a death line is considered, but is clearly rejected for $\dot{P}$ in both case.
 3.2 Simulating 1 millions pulsars
@@
 The parameters used for these runs are shown in Table 5.
-This new set of simulations is intended to verify if generating pulsars younger than $4.1\times 10^{7}$ yr is sufficient, since we observe a small excess of very old pulsars.
+This new set of simulations is intended to verify if generating pulsars younger than $4.1\times 10^{7}$ yr is sufficient, since older pulsars are barely observed.
 The death line is also taken into account.
+Without any specifications, in all figures, with the exception of fig.
+12, 13 and 14, simulations were performed while considering the ISM effect.
 Table 8: Number of pulsars detected with the death line implementation for 1 million pulsars simulated.
 log($\dot{E}$) (W) $N_{tot}$ $N_{r}$ $N_{g}$ $N_{rg}$
-¿31 2 0 2 0
-¿28 124 7 73 44
-Total 1694 1414 154 126
+¿31 8 0 7 1
+¿28 138 6 82 52
+Total 1645 1338 165 142
 Figure 8: Distribution of the observed period taken from the ATNF catalogue, along with the simulations with the death line implementation for 1 million pulsars simulated.
@@
 8 and 9, we conclude that observations and simulations are very similar in both histograms.
-A KS test was conducted for both distributions $P$ and $\dot{P}$ and yielded p-values of $0.24$ and $0.07$ respectively.
-Therefore, it is impossible to reject the null hypothesis that the simulated distributions are similar to the observed distribution.
+A KS test was conducted for both distributions $P$ and $\dot{P}$ and yielded p-values of $0.67$ and $0.17$ respectively.
+Therefore, it is impossible to reject the null hypothesis stating that the simulated distributions are similar to the observed distribution.
 Figure 11: Density plot of the $P-\dot{P}$ diagram in comparison with observations for 1 million pulsars simulated.
 Fig 10 shows the good agreement between simulations and observations.
-Furthermore, we can compare now Fig.
+Furthermore, comparing Fig.
 11 and Fig.
-7.
-It is obvious that Fig.
-11 is better, because globally the colors are lighter which signifies the proportion of pulsars in each 2D bin are almost alike.
-In addition, there is not a pile up effect anymore next to the death line, which corresponds better to the data.
+7, the former performs better because globally the values of $R$ are close to zero, highlighting the fact that the proportion of pulsars in each 2D bin are almost alike.
+In addition, the pile up effect close to the death line has been removed from the $P-\dot{P}$ diagram, which better fits to the data.
 Figure 12: Distance distribution to Earth for observed and simulated populations (1 million pulsars), including the death line implementation.
-The distribution of the distance to Earth is shown in Fig.
+The pulsar distance distribution to Earth is shown in Fig.
 12.
-It has a slightly better similarity to the observations than the work of Dirson et al.
-(2022).
-Nevertheless the improvement is marginal compared to the improvements of $P$ and $\dot{P}$ histograms.
-However, we noticed that removing the effect of dispersion and scattering of the ISM on the radio pulse profile, we would obtain very similar distribution of distances between observations and simulations.
-Therefore, a refinement in the description of the effect of the ISM on the pulse profile might be necessary.
-Indeed, Fig. 13 confirms that the positions of the pulsars in the X-Y plane of the Galaxy are slightly more scattered than before, but the improvement is marginal.
+The improvement compared to Dirson et al.
+(2022) is marginal.
+Anyway, let us remind that the estimated distances of the observed pulsars are subject to large uncertainties because of the distances derived from the dispersion measure values (DM) are highly uncertain, while the distances for the simulated pulsars are correct by construction.
+As a consequence, the comparison between the observed and simulated distance distribution is not a reliable indicator to check the correctness of our model as the discrepancies are subject to observational uncertainties.
+Fortunately the $P-\dot{P}$ improved compared to the previous work.
+This could mean the observed pulsars are closer than expected.
+Indeed, Fig.
+13 shows that the positions of the simulated pulsars in the X-Y plane of the Galaxy, including the ISM effect.
+They are closer to the Sun than the observed pulsars.
+Surprisingly, by removing the effect of dispersion and scattering of the ISM on the radio pulse profile, we obtain very similar distance distributions between observations and simulations, see Fig.
+12.
+Therefore, a refinement of the ISM impact on radio pulse profiles might help getting a more realistic model.
 Figure 13: Spatial distribution of the simulated pulsars, along with the observations, projected onto the Galactic plane for 1 million pulsars simulated.
@@
 Finally fig.
-14 compares the Galactic latitudes from the previous study (see Fig.
-9 of Dirson et al.
-(2022)), now the simulated distribution resembles more the observations, especially around all the latitudes that are different from 0°.
+14 compares the Galactic latitude distributions with the study of Dirson et al.
+(2022), their Fig.
+9.
+In our work the simulated distribution better matches the observations, especially around all the latitudes away from 0°.
 Therefore, the Galactic potential helped tracking the pulsar trajectories more realistically.
+We notice also the ISM removes pulsars close to a latitude of 0°, because the electron density is highest in the galactic plane of the Milky Way.
 3.2.2 Without a death line
-Because the simulation where 1 million pulsars are simulated do not allow to obtain pulsars older than $4.1\times 10^{7}$ yr, we undertook a thorough examination of the usefulness of a death line in modeling the canonical pulsar population.
-Upon conducting simulations with and without a death line (see Table 5 for the parameter values used in the simulations), it emerged that the resultant number of detected pulsars and the corresponding plots exhibited remarkable similarities.
-In addition, when we conduct the KS test on $P$ and $\dot{P}$ between 2 simulations with and without death line, we get p-values of 0.71 and 0.16 respectively, which confirms the similarity between the two simulations.
+Simulation with 1 million simulated pulsars does not allow to generate pulsars older than $4.1\times 10^{7}$ yr.
+We therefore undertook a thorough examination of the usefulness of a death line in modeling the canonical pulsar population.
+Upon conducting simulations with and without a death line (see Table 5 for the parameter values used in the simulations), we found that the resultant number of detected pulsars and the corresponding plots exhibited remarkable similarities.
+By conducting the KS test on $P$ and $\dot{P}$ between two simulations, one with and one without the death line, we obtained p-values of 0.20 and 0.16, respectively.
+Since both p-values are greater than 0.05, this indicates no significant difference between the two simulations, supporting their similarity.
 Consequently, our attention turned towards exploring the impact of lower birth rates to ascertain if they provide a more congruous fit for both spin period and spin period derivative distributions.
@@
 Intriguingly, a discernible paucity of pulsars congregating in the bottom right quadrant of the diagram, colloquially known as the pulsar graveyard, persisted even under these conditions, therefore regardless of the birth rates, thus casting doubt on the necessity of a death line in this configuration.
-Ultimately, our analysis underscores the pivotal importance of delineating the maximum age of simulated pulsars, as elucidated in Subsection 3.1, while the incorporation of a death line do not seem indispensable for attaining realistic results when simulating 1 million pulsars.
+Ultimately, our analysis underscores the pivotal importance of delineating the maximum real age of simulated pulsars, as elucidated in Subsection 3.1, while the incorporation of a death line do not seem indispensable for attaining realistic results when simulating 1 million pulsars.
 3.3 Spin-velocity misalignment effect
@@
 Noutsos et al.
-(2013) showed by studying a catalog of 58 pulsars that pulsars younger than 10 Myr keep their rotation axis and their velocities vector still aligned or anti-aligned, because they were aligned at birth according to Rankin (2007).
+(2013) showed by studying a catalogue of 58 pulsars that pulsars younger than 10 Myr keep their rotation axis and their velocities vector still aligned or anti-aligned, because they were aligned at birth according to Rankin (2007).
 Meanwhile, pulsars older than 10 Myr do not show this trend anymore.
 This progressive misalignment effect is due to the movement through the Galactic potential.
+Fig.
+17 demonstrates that most pulsars have their rotation axis aligned or counter-aligned with their velocities, but for spin-velocity angles between 30° and 150° the numbers of pulsars is considerably lower than around alignment 0° and counter-alignment 180°.
+On the one hand, for pulsars younger than 10 Myr, shown in blue in Fig.
+18, only a small fraction of pulsars have their spin-velocity angle greater than 30° and below 150°.
+On the other hand, for pulsars older than 10 Myr, shown in red in Fig.
+18, their spin-velocity angle becomes greater than 30° or smaller than 150°.
+Furthermore, Fig.
+19 confirms that, the older the pulsar is, the better its chance to perform several orbits in the Galaxy, allowing it to significantly deviate from spin-velocity alignment or anti-alignment.
+In addition, even for less old pulsars able to perform many orbits in the Galaxy, their spin-velocity angle will evolve between 30° and 150°.
+This supports the idea that the Galactic potential is responsible for the misalignment irrespective of the age of the pulsar.
+The opposite is also true, some old pulsars do not perform many orbits (but most of them are), and as a consequence they keep approximately their spin-velocity alignment or misalignment.
+The previous plots clearly highlight the influence of the Galactic potential making pulsars lose their alignment between their rotation axis and their proper motion velocities through time.
 Figure 17: Distribution of the spin-velocity angles for all the detected pulsars in the simulation for 1 million pulsars simulated.
 Figure 18: Distribution of the spin-velocity angles for the detected pulsars younger than 10 Myr (blue) and pulsars older than 10 Myr (red) in the simulation.
-Fig.
-17 demonstrates that most pulsars have their rotation axis aligned with their velocities, but for spin-velocity angles between 25° and 180° the numbers of pulsars is considerably lower than around 0° and 180°.
-On the one hand, for pulsars younger than 10 Myr, shown in blue in Fig.
-18, only a small fraction of pulsars have their spin-velocity angle greater than 25° and below 150°.
-On the other hand, for pulsars older than 10 Myr, shown in red in Fig.
-18, their spin-velocity angle becomes greater than 25° or below 150°.
 Figure 19: The spin-velocity angles for the detected pulsars in function of the number of orbits made by the pulsars.
-Furthermore, in Fig.
-19, we can confirm that, the older the pulsar is, the better is its chance to orbit more in the Galaxy, allowing it to deviate its spin-velocity angle from alignment (or anti-alignment).
-In addition, even if the pulsar is not that old, if it made many orbits in the Galaxy, its spin-velocity angle will be between 25° and 150°, which confirms the fact that its the Galactic potential (and not only age) which makes this spin-velocity angle increase (or decrease if its is anti-aligned at birth).
-In the contrary, there are pulsars which are old which did not made many orbits in the Galaxy, as a consequence they have their spin-velocity angle near 0° or 180°.
-These plots clearly shows the influence of the Galactic potential making the pulsars lose their alignment between their rotation axis and their velocities through time because they evolve in the Galactic potential.
 3.4 Log normal vs normal distribution for spin period at birth
-Figure 20: Distribution of the observed period taken from the ATNF catalogue, along with the simulations with the death line implementation and the Gaussian distribution at birth for the spin period.
-Table 9: Number of pulsars detected with the death line implementation and the Gaussian distribution at birth for the spin period.
-log($\dot{E}$) (W) $N_{tot}$ $N_{r}$ $N_{g}$ $N_{rg}$
-¿31 7 0 5 2
-¿28 275 9 138 128
-Total 1642 1188 223 231
-In order to estimate the impact of a new distribution of spin period at birth, we compare the results with the previous distribution of spin period
-$p(P_{0})=\frac{1}{\sigma_{p}\sqrt{2\pi}}e^{-(P_{0}-\bar{P})^{2}/(2\sigma_{p})^{2}}$ (34)
-The simulation in this subsection uses $\bar{P}=60$ ms and $\sigma_{p}=10$ ms as parameters of equation (34).
-The Galactic potential and the death line were both taken into account to compare the influence of the initial spin period distribution and 1 million pulsars were simulated, see Table 5 for the parameters used.
-Table 9 highlights the fact that much more pulsars are detected above $\dot{E}>10^{28}$ W compared to the simulation with the Gaussian in log space for the spin period, see Table 7.
-It confirms the suspicion that the distribution of initial spin period influences the number of pulsars at higher spin down luminosity.
-The spin period distribution in Fig.
-20 shows some simulated pulsars with a low spin period, in excess compared to the observations.
-This artifact disappears when the Gaussian distribution in log space for the spin period is used.
-The log-normal distribution enables to remove those pulsars in excess and make the distribution of the spin period better fit the observations.
-The KS test was also conducted on both $P$ and $\dot{P}$ and we find p-values, once again, considerably below 0.05 for both quantity, therefore rejecting the null hypothesis.
+Table 9: P-values and total number of pulsars detected ($N_{detection}$) obtained for simulations with different $\bar{P}$ for the Gaussian initial distribution of spin period.
+$\bar{P}$ (ms) $\sigma_{p}$ (ms) p-value (P) p-value ($\dot{P}$) $N_{detection}$
+60 30 $6.58\times 10^{-20}$ $3.76\times 10^{-4}$ 1512
+80 30 $1.84\times 10^{-15}$ 0.051 1671
+100 30 $2.34\times 10^{-10}$ 0.123 1721
+129 30 $1.02\times 10^{-6}$ 0.050 1787
+140 30 $1.83\times 10^{-5}$ 0.006 1816
+In order to estimate the impact of the spin period distribution at birth, we compare the results of the log-normal law with the previously used normal law in Dirson et al. (2022)
+$p(P_{0})=\frac{1}{\sigma_{p}\sqrt{2\pi}}e^{-(P_{0}-\bar{P})^{2}/(2\sigma_{p})^{2}}\ .$ (36)
+The simulation in this subsection uses the parameters listed in the two first columns of Table 9 and to be put into equation (36).
+The Galactic potential and the death line were both taken into account to compare the influence of the initial spin period distribution on the 1 million pulsars simulated, see Table 5 for the parameters used.
+In Table 9, only results for $\sigma_{p}=30$ ms are shown because other values have been tried out, namely 10 ms, 20 ms and 40 ms but gave no higher p-values than for $\sigma_{p}=30$ ms.
+Moreover, we chose $\bar{P}$ from 60 ms to 140 ms because it is the range usually used in other work (Faucher-Giguère & Kaspi, 2006; Gullon et al., 2014; Dirson et al., 2022).
+Table 9 highlights the fact that for $\bar{P}$ above 60 ms and below 140 ms, an acceptable p-value $\geq 0.05$ is reached for the $\dot{P}$ distribution.
+However no matter the value of $\bar{P}$, a satisfying p-value for the $P$ distribution is not reached in contrary to the log-normal distribution at birth, for which satisfying p-value for both distributions are obtained.
 We conclude that the initial distribution for the spin period found by Igoshev et al.
 (2022) allows to better reproduce the canonical pulsars population.
-3.5 Gamma-ray detection
-Let us now focus on the $\gamma$-ray band of the pulsar population.
-First, we note that the total number of pulsars detected simultaneously in gamma-ray and in radio is significantly larger in our simulation compared to the observations, when using the Fermi/LAT instrument sensitivity.
-However, the simulated gamma-ray pulsars are in the expected area of Fig.
-21, meaning the striped wind model allow to correctly reproduce these pulsars.
-The explanation of this higher number of detection would be that there are many unidentified sources in the Fourth Fermi-LAT Catalog of Gamma-ray sources (4FGL).
-Since many unidentified Galactic sources are listed in these catalogs, many gamma-ray pulsars might have been detected without being identified as such yet.
-Furthermore, we can compare our results with the PPS work of Pierbattista et al.
-(2012), where they simulated the young gamma-ray pulsars population with 4 different models : Outer Gap (OG), Slot Gap (SG), Polar Cap (PC) and Outer Polar Cap (OPC) models.
-In summary their best model was the OPC model, which shows almost identical results as ours (the $P-\dot{P}$ shape and the number of detection), except the fact that we detect gamma-ray pulsars a little further away from Earth.
-Figure 21: $P-\dot{P}$ diagram of the gamma-only pulsars for both the simulations and the observations, with 1 million pulsars simulated.
-What if we run another simulation (with the same parameters as in Table 5), by only changing the threshold of detection for gamma-ray ?
-We did this in order to see how many more pulsars we would detect in gamma-ray and where they would be in the $P-\dot{P}$ diagram, and also to deal with the problematic of the GeV excess in the Galactic centre.
-Therefore we increased $F_{\rm min}$ by 10 : if the galactic latitudes of the pulsar is ¡ 2°, then $F_{\rm min}=0.4\times 10^{-15}$ W.m-2, and concerning blind searches $F_{\rm min}=1.6\times 10^{-15}$ W.m-2.
-Figure 22: Spatial distribution of the simulated gamma-ray detected pulsars, projected onto the Galactic plane for 1 million pulsars simulated, with a lower $F_{min}$.
-The gamma-ray detected pulsar positions in the X-Y plane of the Galaxy are shown in Fig.
+3.5 $\gamma$-ray detection
+Let us now focus on the $\gamma$-ray pulsar population.
+First, we note that the total number of pulsars detected is significantly larger in our sample (307 detections) compared to the observations (173 detections), when using the Fermi/LAT instrument sensitivity.
+However, the simulated $\gamma$-ray only and radio loud $\gamma$-ray pulsars are in the expected area of Fig.
+20, furthermore, the KS test gives a p-value of 0.52 and $\approx 0$ for $\dot{P}$ and $P$ respectively.
+The null hypothesis can not be rejected for $\dot{P}$ but it is rejected for $P$.
+However, when the KS test is conducted on the $\gamma$-ray population only we get p-values of 0.03 and 0.14 for $\dot{P}$ and $P$ respectively, therefore here the null hypothesis is rejected for $\dot{P}$ but not for $P$.
+While for the radio loud $\gamma$-ray population p-values of 0.11 and $\approx 0$ are obtained for $\dot{P}$ and $P$ respectively.
+Thus, we thought the low p-value obtained for $P$ for both population was probably because of the 7 high spin period radio loud $\gamma$-ray pulsars seen on Fig.
+20, but even when conducting the KS test without these pulsars gave a p-value $\geq 0.05$ for $\dot{P}$ and well below $0.05$ for $P$.
+In addition, we also compare the $\gamma$-ray flux distribution of the simulated and observed populations.
+In Fig.21 we notice that both distributions peak at the same flux level around $10^{-13.7}$ W.m-2.
+Globally, the number of pulsars in the simulation with a flux greater than $10^{-13.4}$ W.m-2 is very close to the observations.
+However, between $10^{-14.4}$ W.m-2 and $10^{-14}$ W.m-2, there are much more pulsars in the simulation.
+This bi-modality in the simulated $\gamma$ray flux distribution is an artefact introduced by our abrupt condition on the threshold $F_{\rm min}$ depending on the radio detection of a pulsar and on a sufficiently low galactic latitude.
+However if we smoothly change $F_{\rm min}$ from $4\times 10^{-15}$ to $16\times 10^{-15}$ (the two $F_{\rm min}$ values used in this work), the bi-modality would almost disappear.
+The higher number of detected pulsars in the simulations could be related to the unidentified sources in the Fourth Fermi-LAT catalogue of $\gamma$-ray sources (4FGL).
+Since many unidentified Galactic sources are listed in this catalogue, many $\gamma$-ray pulsars might have been observed without being identified as such yet.
+Nonetheless, assuming that every unidentified $\gamma$-ray pulsar in this catalogue will be recognized as such by a future more sensitive instrument and that a significant difference between the number of simulated and observed pulsars subsists, this could hint to some shortcomings faced by the striped wind model.
+Let us mention some of these possible flaws of geometrical or physical nature.
+First the $\gamma$-ray beam opening angle could be smaller than the one used due to some plasma recollimation effect within the striped wind.
+This would reduce the number of detected pulsars.
+Second the luminosity function could also deviate from the prescription adopted in (32).
+Another dependence on $B$ and $P$ or equivalently on $P$ and $\dot{P}$ could be introduced, maybe adding a third parameter like the cut-off energy to better describe the $\gamma$-ray luminosity.
+Furthermore, we compare our results with the PPS work of Pierbattista et al.
+(2012), where they simulated the young $\gamma$-ray pulsar population with 4 different models : Outer Gap (OG), Slot Gap (SG), Polar Cap (PC) and Outer Polar Cap (OPC) models.
+To summarize their findings, their best model is the OPC model, showing almost identical results as ours, $P-\dot{P}$ shape and number of detection, except that we detect $\gamma$-ray pulsars a little farther away from Earth.
+Figure 20: $P-\dot{P}$ diagram of the $\gamma$-only pulsars for both the simulations and the observations, with 1 million pulsars simulated.
+Figure 21: Distribution of the $\gamma$-ray flux of the simulated population along with the observations from the 3PC catalogue.
+To conclude the part about $\gamma$-ray pulsars, we predict the number of pulsars that would be detected by a future instrument, ten times more sensitive than Fermi/LAT.
+We asked the question: what if we run another simulation with the same parameters as in Table 5, only changing the threshold of detection for $\gamma$-ray?
+This investigation could help in identifying the nature of the GeV excess in the Galactic centre.
+Therefore we decreased the instrument sensitivity $F_{\rm min}$ by a factor 10: if the galactic latitudes of the pulsar is ¡ 2°, then $F_{\rm min}=0.4\times 10^{-15}$ W.m-2, and concerning blind searches we set $F_{\rm min}=1.6\times 10^{-15}$ W.m-2.
+Figure 22: Spatial distribution of the simulated $\gamma$-ray detected pulsars, projected onto the Galactic plane for 1 million pulsars simulated, with a lower $F_{min}$.
+The $\gamma$-ray detected pulsar positions in the X-Y plane of the Galaxy are shown in Fig.
 22.
-In this simulation, 1832 gamma-ray pulsars are detected, among them only 168, in red in Fig.
-22, would have been detected if the same sensitivity as before was kept.
-Therefore we detect 11 times more gamma-ray pulsars compared to the number found with the previous sensitivity.
-Furthermore, the other 1664 pulsars, in green in Fig.
-22, detected with this increased sensitivity highlight several features.
-We obviously detect more pulsars close to us, and also many more in the centre of the Galaxy.
-This result indicates that the GeV excess in the Galactic centre could be linked to pulsars not yet identified, while usually the hypothesis is that the GeV excess corresponds to self-annihilating dark matter particles (Hooper, 2022).
-In addition, millisecond pulsars, not modeled in this work, could however also contribute to this GeV excess.
+In this simulation, 2492 $\gamma$-ray pulsars are detected, in green, being $\gamma$-ray only or radio-loud $\gamma$-ray pulsars, among them only 350 (really close to the number of $\gamma$-ray and radio/$\gamma$-ray pulsars detected in Subsection 3.2) would have been detected if the same sensitivity as before was kept, in red in Fig.
+22.
+Therefore we detect 7 times more $\gamma$-ray pulsars compared to the number found with the previous sensitivity.
+Furthermore, the other 2142 pulsars, in green in Fig.
+22, detected with this increased sensitivity highlight several interesting features.
+We obviously detect more pulsars close to Earth, and also many more in the centre of the Galaxy.
+This result indicates that the GeV excess in the Galactic centre could be linked to pulsars not yet identified, while usually the origin of this GeV excess is attributed to self-annihilating dark matter particles (Hooper, 2022).
+However, millisecond pulsars, not modeled in this work, could also contribute significantly to this GeV excess.
 4 Discussion
 For the initial spin period distribution Igoshev et al.
-(2022) suggested values for the mean period $P_{\text{mean}}$ in the range from approximately $57$ ms to $129$ ms and for the spread $\sigma_{p}$ the range from $0.45$ to $0.65$.
+(2022) suggested values for the mean period $P_{\text{mean}}$ in the range from approximately $57$ ms to $129$ ms and for its spread $\sigma_{p}$ the range from $0.45$ to $0.65$.
 The values $P_{\text{mean}}=129$ ms and $\sigma_{p}=0.45$ are best educated guesses.
-However it is obvious that further investigations could benefit from an automated method to find the best set of values.
-Hence, we are currently working on an optimization algorithm to find the optimal initial parameters for the population synthesis, not only for the spin period, but for the whole parameter space presented in Table 5, except for $\alpha_{d}$ and $k_{\tau_{d}}$.
+However further investigations on this work could benefit from an automated method to find the best set of values.
+Hence, we are currently working on an optimization algorithm to find the optimal initial parameters for the population synthesis, not only for the spin period, but for the whole parameter space presented in Table 5.
 Our optimization algorithm will help to constrain the properties of the Milky Way pulsar population by constraining parameters like the birth rate which is supposed to be between $1/33$ yr-1 and $1/150$ yr-1, see Faucher-Giguère & Kaspi (2006), Gullon et al.
 (2014) or Johnston & Karastergiou (2017).
-We also added the effect of the ISM on the radio pulse profile, rendering more difficult to observe pulsars farther away than $10$ kpc.
-Furthermore, because of this plasma propagation effect fewer radio pulsars are detected compared to the observations.
-If the simulations discard the ISM effect, much more pulsars are detected at distances larger than $10$ kpc and with higher $\dot{E}$ as well, improving the match between the simulation and the observation.
-Perhaps, the effect of the ISM is overestimated when computing the scattering and dispersion effect converted in equation (28a) with $\tau_{scat}$ and $\tau_{DM}$ is too important and should have less influence on the observed width pulse profile, therefore with formulas that would give lower $\tau_{DM}$ and $\tau_{scat}$, we might get more realistic results.
 In the main part of the paper, we abstain from showing outcomes stemming from simulations featuring a constant magnetic field.
@@
 Moreover appendix C explores different parameter spaces for the PPS in order to meticulously scrutinizing whether the parameters showcased in the main part of the paper optimally encapsulate observational data.
-This demonstrates that even if a substantial number of parameters are required for the PPS, they can be meaningfully constrained by such studies with reasonable accuracy.
-Furthermore, in Subsection 3.5 we explore the results when the sensitivity in gamma-ray is increased by a factor of 10.
-However we did not show what would happen if we tried to make a prediction for the radio detection by increasing the sensitivity by a factor of 50 to match approximately the properties of SKA.
-A simulation was run with this increased sensitivity and 52276 pulsars were detected (when simulating 1 million pulsars, with a death line), therefore approximately 31 times more than with the sensitivity of Parkes and Arecibo telescopes.
+This work demonstrates that even if a substantial number of parameters are required for the PPS, they can be meaningfully constrained by such studies with reasonable accuracy.
+Furthermore, in Subsection 3.5 we explored the results when the sensitivity in $\gamma$-ray is increased by a factor of 10.
+However we did not discuss improvements in the radio detection yet, while we run a simulation by increasing the sensitivity by a factor of 50 to match approximately the properties of SKA (McMullin et al., 2020).
+The result of this simulation with an increased sensitivity led to 52276 pulsars detected (when simulating 1 million pulsars, with a death line), therefore approximately 31 times more than with the sensitivity of Parkes and Arecibo telescopes.
+We emphasize that the death valley strongly depends on the chosen model.
+For instance the radio luminosity law plays a central role, we assumed that $F_{r}\propto\dot{E}^{\alpha}$ with $\alpha=1/4$ (see equation (27)) according to Johnston & Karastergiou (2017) and Johnston et al.
+(2020) where they found it to best fit to the data.
+Following their conclusion we decided not to choose $\alpha$ as another parameter.
+However it would impact the death valley, but this effect is not considered in our work.
+Furthermore, in other works like Graber et al.
+(2024) for instance, they do not explicitly use a death valley because they use a different radio luminosity law, meaning a different value for $\alpha$ and adjust the magnetic field decay time accordingly.
+We stress that this is just another method to deal with the bottom right corner of the $P-\dot{P}$ diagram, where pulsars are likely to stop emitting photons, since the death valley removes pulsars based on a combination of $P$ and $\dot{P}$.
+Graber et al.
+(2024) PPS is based on a phenomenological approach where the radio luminosity is tuned to conform to observations.
+In our approach, adding a death valley is based on a more physical ground related to the microphysics close to the neutron star polar cap.
 5 Summary
@@
 This evolution of the proper motion was not realistic, even though the evolution model had already replicated a significant portion of the canonical pulsar population.
-However all the new improvements put into this PPS, meaning : the Galactic potential, the effect of the ISM on the radio pulse profiles and the fact that the decay timescale of the magnetic field is taken randomly within a certain range, allow a more accurate reproduction of the population in the $P-\dot{P}$ diagram.
+However all the new improvements put into this PPS, meaning the Galactic potential, the effect of the ISM on the radio pulse profiles and the fact that the decay timescale of the magnetic field is taken randomly within a certain range, allow a more accurate reproduction of the population in the $P-\dot{P}$ diagram.
 The first results where we emulate 10 millions pulsars showed an excess of old pulsars detected in the simulation.
@@
 The results obtained with the death line were then in better agreement with observations when plotting the most relevant distributions such as period and period derivative.
-Nonetheless, simulating 1 million pulsars allowed to show that with or without death line the results were alike and therefore, in the case where the oldest pulsar simulated is $4.1\times 10^{7}$ yr, a death line is not really necessary to explain the observations.
-In this paper we showed two approach to reproduce the observed canonical pulsars population, however when we consider the approach where the oldest pulsar simulated is only $4.1\times 10^{7}$ yr we state that no older pulsars might exist in the Galaxy, while the Milky Way is very much older than this.
-There is apparently no reason to not simulate pulsars that are older than this, and if there is one we do not know it (they become harder to detect because of their wide pulse profile with their age, but some are still detectable).
-If this situation is the reality then it means two things : the age of the observed pulsars are overestimated (because we observe pulsars older than $4.1\times 10^{7}$ yr in the canonical population, even if they are not numerous) and the death line is not necessary in pulsar population synthesis work to explain what we observe.
+Nonetheless, simulating 1 million pulsars allowed to show that with or without a death line the results were alike and therefore, in the case where the oldest simulated pulsar is $4.1\times 10^{7}$ yr (its real age), a death line is not necessary to explain the observations.
+In this paper we showed two approaches to reproduce the observed canonical pulsar population, however when we consider the approach where the oldest pulsar simulated is only $4.1\times 10^{7}$ yr we state that no older pulsars might exist in the Galaxy, while the Milky Way is very much older than this.
+There is apparently no reason not to simulate pulsars that are older than this, and if there is one we do not know it, they become harder to detect because of their wide pulse profile with their age, but some are still detectable.
+If this situation is the reality then it means that the death line is not necessary in pulsar population synthesis to explain observations.
 When the simulation is ran with 10 millions pulsars, the death line becomes a necessity.
-However this simulation gives results less similar to the observations, because of a few simulated data which do not match any observational data.
-It was also shown that the old pulsars with age larger than 10 Myr have a tendency to lose their alignment between rotation axis and velocity, a direct effect of the Galactic potential as shown by the observations of Noutsos et al.
+However this simulation gives results less similar to observations.
+It was also shown that old pulsars with age larger than 10 Myr have a tendency to lose their alignment between rotation axis and velocity, a direct effect of the Galactic potential as shown by the observations of Noutsos et al.
 (2013), but it is the first time it is done in a simulation.
-The striped wind model seems to reproduce the gamma-ray population of canonical pulsars very well, however in the simulation a higher number of these pulsars are detected compared to reality.
-It probably indicates that a part of unidentified sources from the 4FGL catalog are pulsars.
-Testing an increased sensitivity for the detection of gamma-ray pulsars lead us to think that a part of the origin of the GeV excess in the Galactic centre could be young gamma-ray pulsars, and when the millisecond pulsars will be modeled in our simulation we would also be able to check that these pulsars might be responsible for this GeV excess.
-The usual picture is to suppose that only millisecond pulsars are responsible, while young gamma-ray pulsars could also play a role.
-In summary, the best parameters found for this pulsar population synthesis are for $P_{\text{mean}}=129$ ms, $\sigma_{p}=0.45$ for the spin period distribution at birth, $B_{\text{mean}}=2.75\times 10^{8}$ T, $\sigma_{b}=0.5$ for the magnetic field distribution at birth and for a birth rate of 1/41 yr-1.
+The striped wind model seems to reproduce the $\gamma$-ray population of canonical pulsars well according to the p-value of the KS test greater than 0.05 for $\dot{P}$.
+However the KS test gives a p-value lower than 0.05 for $P$, but there are several observed radio-loud $\gamma$-ray pulsars that the simulation struggles to reproduce (not only the ones with high $P$) and might explain this low p-value.
+However in the simulation a higher number of these pulsars are detected compared to observations.
+It probably indicates that a part of the unidentified sources from the 4FGL catalogue are pulsars.
+Testing an instruments with improved sensitivity for the detection of $\gamma$-ray pulsars lead us to the conclusion that part of the GeV excess in the Galactic centre could be young $\gamma$-ray pulsars.
+Moreover the assumption that millisecond pulsars could also contribute to this excess will be checked once a PPS model for millisecond pulsars will be available.
+The most common picture suppose that only millisecond pulsars are responsible for this excess, while we demonstrated that young $\gamma$-ray pulsars could also play a significant role.
+In summary, the best parameters found for this pulsar population synthesis are for $P_{\text{mean}}=129$ ms, $\sigma_{p}=0.45$ for the log-normal spin period distribution at birth, $B_{\text{mean}}=2.75\times 10^{8}$ T, $\sigma_{b}=0.5$ for the magnetic field distribution at birth and for a birth rate of 1/41 yr-1.
 Ultimately, this work provides a simulated population of pulsars within the Milky Way more similar to the observations compared to Johnston et al.
 (2020) and Dirson et al.
-(2022), the two most recent population synthesis considering both radio and gamma-ray emission.
-Moreover, such investigations will be useful to make more predictions for future radio surveys such as SKA, or NenuFAR.
-The PPS method used in this study could be applied to model populations of pulsars in the Milky Way other than the canonical ones.
-The magnetars, for instance, would need : another distribution at birth for their magnetic field to be modeled correctly, taking into account the emission in X-rays and having a faster decay for the magnetic field.
-The magnetars are the pulsars just above the canonical pulsars in the $P-\dot{P}$ diagram, at the top right of the diagram.
+(2022), the two most recent population synthesis considering both radio and $\gamma$-ray emission.
+Moreover, such investigations are useful to predict the detection rate of future radio surveys such as SKA or NenuFAR.
+The PPS method developed in this paper can be extended to other pulsar populations in the Milky Way such as millisecond pulsars and magnetars.
+The magnetars, for instance, would need another distribution at birth for their magnetic field, taking into account the emission in X-rays and having a faster decay for the magnetic field.
+The magnetars are located just above the canonical pulsars in the $P-\dot{P}$ diagram, at the top right of the diagram.
 A more complex population is the recycled or millisecond pulsars, that were ”brought back to life” by accreting matter from a companion star.
-This accretion mechanism spins up the neutron stars moving them from the graveyard of pulsars to the bottom left of the $P-\dot{P}$ diagram, crossing again the death line in the opposite sense and allowing them to create electron-positron pair cascades again.
-However, they are more complex to model because this evolution scenario should take into account the accretion process and spin-up.
-We are also currently working on both populations to reproduce them in a PPS based on this work.
+This accretion mechanism spins up the neutron stars moving them from the graveyard of pulsars to the bottom left of the $P-\dot{P}$ diagram, crossing again the death line in the opposite sense and allowing them to revive by creating electron-positron pair cascades again.
+However, they are difficult to model because their evolution scenario must take into account their accretion phase and spin-up.
+We currently work on both these populations to reproduce them in a PPS based on this work.
 Acknowledgements.
@@
 here ${r}$ is the position vector, ${v}$ is the velocity vector, $h$ is the time step, ${\nabla}\Phi$ is the gradient of the potential and $n$ is the time at which we compute the position or the velocity.
-$\lambda$, $\xi$ and $\chi$ are constants that are tabulated in Table 10, see for instance Peter Young (2014).
+$\lambda$, $\xi$ and $\chi$ are constants that are tabulated in Table 10, see for instance Young (2014).
 Table 10: Value of the constants used for the PEFRL scheme.
@@
 Hence, this simulation was run with 1 million pulsar simulated, without a death line and without a decaying of the magnetic field.
-The KS test gave p-values of $6.3\times 10^{-22}$ and $9.8\times 10^{-77}$, therefore the null hypothesis is rejected for both $P$ and $\dot{P}$.
+The KS test gave p-values of $1.1\times 10^{-14}$ and $1.5\times 10^{-76}$, therefore the null hypothesis is rejected for both $P$ and $\dot{P}$.
 The parameters used are in Table 5.
-It is less good than with a decaying magnetic field.
+It gives results less close to the observations than with a decaying magnetic field.
 As can be seen in Fig.
@@
 3.
+We could expect to reproduce equally well the population of observed pulsars with a constant magnetic field, however compared to the study of Johnston & Karastergiou (2017) where they used this assumption for the magnetic field, we do not evolve the braking index.
+In our simulations, we have $\dot{\alpha}<0$, $\dot{B}<0$ and $\dot{n}=0$ whereas in their simulation $\dot{\alpha}<0$, $\dot{B}=0$ and $\dot{n}\neq 0$.
+Having either the braking index or the magnetic field evolve, allows to reproduce well the population of pulsars.
 Figure 25: $P-\dot{P}$ diagram of the simulated population, along with the observations, with the death line implementation and a constant magnetic field for each pulsar.
@@
 log($\dot{E}$) (W) $N_{tot}$ $N_{r}$ $N_{g}$ $N_{rg}$
-¿31 15 0 12 3
-¿28 168 4 94 70
-Total 1356 1034 162 160
+¿31 14 0 10 4
+¿28 157 6 77 74
+Total 1328 974 176 178
 Appendix C Influence of the different parameters
-In order to show that the parameters chosen in the main part of this paper are meticulously picked and that the uncertainties brought by the large number of parameters is manageable, several simulations were run with different values for the main parameters of the initial distributions, showing a deviation from the best values gives bad results.
-The simulations were run with in the same conditions as Subsection 3.2, with only one parameter changing between the different sets of simulation.
-Table 12: Influence of the birth rate.
-Birth rate (in yr-1) $N_{tot}$ $N_{r}$ $N_{g}$ $N_{rg}$
-1/30 2388 1964 204 220
-1/80 963 799 79 85
-1/120 696 597 50 49
-Table 13: Influence of the mean initial period.
-$P_{\rm mean}$ (in ms) $N_{tot}$ $N_{r}$ $N_{g}$ $N_{rg}$
-57 1499 1134 193 172
-70 1623 1253 215 155
-100 1652 1348 151 153
-Table 14: Influence of the mean initial magnetic field.
-$B_{mean}$ (in T) $N_{tot}$ $N_{r}$ $N_{g}$ $N_{rg}$
-$10^{7}$ 8407 6943 406 1058
-$5\times 10^{7}$ 4447 3646 306 495
-$5\times 10^{8}$ $1$ 933 109 69
-The $P-\dot{P}$ diagram of each simulation is not shown here because almost every simulation finds a too high (or too low) number of pulsars, therefore the $P-\dot{P}$ diagrams are showing large differences between observations and simulations.
-Nonetheless, the number of pulsars detected for two of the simulations with different $P_{\rm mean}$ (with $P_{\rm mean}=70$ ms and $P_{\rm mean}=100$ ms) gives number closer to reality.
-However their p-values for $P$ and $\dot{P}$ when using the KS test were significantly below $0.05$, which allowed us to reject the null hypothesis for the distributions of $P$ and $\dot{P}$, for the results shown in Table 12, 13 and 14.
-Moreover in order to get an idea on a good range of values for the parameters, the d-value obtained thanks to the KS test are shown in Table 15, 16, 17 and 18.
-For each parameter, 10 simulations were made in order to get a mean and a standard deviation for the d-value associated.
-These simulations helped finding the best values for the parameters by considering the parameter is well constrained when its d-values for both distribution of $P$ and $\dot{P}$ are below, or really close to 0.05 (meaning 5 % similarity between the observations and simulation distribution).
-Even though it is not an optimisation (and therefore the results can be improved on that point), it allows us to conclude that $P_{\rm mean}$ should be between 120 ms and 135 ms (with 129 ms as best value).
-$B_{\rm mean}$ should be between $2.25\times 10^{8}$ T and $2.8\times 10^{8}$ T (with $2.75\times 10^{8}$ T as best value).
-The birth rate should be between 1/37 yr-1 and 1/45 yr-1 (with 1/41 yr-1 as best value).
-Finally, $\sigma_{b}$ is well constrained between 0.49 and 0.51 (with 0.5 as best value).
-Concerning $\sigma_{p}$, it always gives d-values between 0.04 and 0.057 when $\sigma_{p}$ is below 0.6, therefore the value 0.45 was chosen as suggested by Igoshev et al.
-(2022).
+This appendix shows that the parameters chosen in the main part of this paper are meticulously picked and that the uncertainties brought by the large number of parameters is manageable.
+Multiple simulations were conducted by varying parameter values for the initial distributions, demonstrating that deviations from the optimal values lead to poor results.
+The simulations were run with in the same conditions as subsection 3.2, with only one parameter changing between the different sets of simulations and without implementing the ISM, because it would be too time consuming to perform all these simulations.
+As a consequence in the tables from 12 to 16), the p-values are low, while when including the ISM effect they become larger.
+Therefore, the results and conclusions presented in this section are based on this simplification, which was primarily used to guide the choice of parameters.
+While these assumptions have helped to identify efficient parameters, it is important to stress that the proposed values are not necessarily optimal.
+They serve primarily as a foundation for refining the parameters under more realistic conditions.
+That is why before validating a parameter we checked that the p-values were satisfactory when including the ISM.
+We found each time a higher p-values.
+In order to get an idea on a good range of values for the parameters, the d-values and p-values obtained thanks to the KS test are shown in Table 12, 13, 14, 15 and 16.
+For each parameter, 10 simulations were performed in order to get a mean and a standard deviation for the d-values and p-values associated.
+These simulations helped finding the best values for the parameters by considering the parameter is well constrained when its d-values for both distribution of $P$ and $\dot{P}$ are below or close to 0.05 (meaning 5% similarity between the observed and simulated distributions), especially if the previous and next parameter values have higher d-values and lower p-values.
+As usual p-values above 0.05 means that the null hypothesis can not be rejected, however as explained above, without the ISM effect taken into account the p-values for the $\dot{P}$ distribution do not get too high here.
+Even though it is not an optimisation (and therefore the results can be improved on that point), it allows us to conclude that $P_{\rm mean}$ must be between 120 ms and 135 ms, with 129 ms as best value.
+$B_{\rm mean}$ must be between $2.25\times 10^{8}$ T and $2.8\times 10^{8}$ T, with $2.75\times 10^{8}$ T as best value.
+$\sigma_{b}$ is well constrained between 0.49 and 0.51, with 0.5 as best value.
+While $\sigma_{p}$ must be between 0.42 and 0.47, with 0.45 as best value.
+Finally, the birth rate must be between 1/37 yr-1 and 1/45 yr-1, with 1/41 yr-1 as best value.
 Table 12: D-value for different $P_{\rm mean}$.
-$P_{\rm mean}$ (in ms) d-value ($\dot{P}$) d-value ($P$)
-90 $0.072\pm 0.0073$ $0.0577\pm 0.0080$
-100 $0.0651\pm 0.00692$ $0.0503\pm 0.00636$
-110 $0.0652\pm 0.00688$ $0.0381\pm 0.0074$
-120 $0.0554\pm 0.00714$ $0.0302\pm 0.00468$
-129 $0.0557\pm 0.00740$ $0.0284\pm 0.00583$
-135 $0.0544\pm 0.00768$ $0.0329\pm 0.00767$
-150 $0.0572\pm 0.00779$ $0.0388\pm 0.00935$
+$P_{\rm mean}$ (in ms) d-value ($\dot{P}$) d-value ($P$) p-value ($\dot{P}$) p-value ($P$)
+90 $0.072\pm 0.0073$ $0.058\pm 0.0080$ $0.0091\pm 0.021$ $0.038\pm 0.049$
+100 $0.065\pm 0.0069$ $0.050\pm 0.0064$ $0.012\pm 0.0089$ $0.22\pm 0.0086$
+110 $0.065\pm 0.0069$ $0.038\pm 0.0074$ $0.0033\pm 0.0022$ $0.32\pm 0.0076$
+120 $0.055\pm 0.0071$ $0.030\pm 0.0047$ $0.0049\pm 0.0060$ $0.47\pm 0.25$
+129 $0.056\pm 0.0074$ $0.028\pm 0.0058$ $0.010\pm 0.014$ $0.22\pm 0.19$
+135 $0.054\pm 0.0077$ $0.033\pm 0.0077$ $0.012\pm 0.015$ $0.091\pm 0.0095$
+150 $0.057\pm 0.0078$ $0.039\pm 0.0093$ $0.011\pm 0.024$ $0.013\pm 0.0074$
 Table 13: D-value for different $B_{\rm mean}$.
-$B_{\rm mean}$ (in T) d-value ($\dot{P}$) d-value ($P$)
- $1\times 10^{8}$ $0.113\pm 0.00823$ $0.203\pm 0.00675$
-$2\times 10^{8}$ $0.0418\pm 0.00459$ $0.0658\pm 0.00705$
-$2.25\times 10^{8}$ $0.0456\pm 0.00455$ $0.0455\pm 0.0061$
-$2.75\times 10^{8}$ $0.0557\pm 0.00740$ $0.0284\pm 0.00583$
-$2.8\times 10^{8}$ $0.0576\pm 0.00907$ $0.0298\pm 0.00728$
-$3\times 10^{8}$ $0.0648\pm 0.01092$ $0.0426\pm 0.00775$
-$3.25\times 10^{8}$ $0.0761\pm 0.00930$ $0.0480\pm 0.00997$
+$B_{\rm mean}$ (in T) d-value ($\dot{P}$) d-value ($P$) p-value ($\dot{P}$) p-value ($P$)
+ $1\times 10^{8}$ $0.11\pm 0.0082$ $0.20\pm 0.0067$ $4.4\times 10^{-16}\pm 1.2\times 10^{-15}$ $8.9\times 10^{-40}\pm 2.4\times 10^{-39}$
+$2\times 10^{8}$ $0.042\pm 0.0046$ $0.066\pm 0.0070$ $0.033\pm 0.043$ $0.0058\pm 0.0072$
+$2.25\times 10^{8}$ $0.046\pm 0.0045$ $0.045\pm 0.0061$ $0.011\pm 0.019$ $0.12\pm 0.076$
+$2.75\times 10^{8}$ $0.056\pm 0.0074$ $0.028\pm 0.0058$ $0.010\pm 0.014$ $0.22\pm 0.19$
+$2.8\times 10^{8}$ $0.058\pm 0.0091$ $0.030\pm 0.0073$ $0.0074\pm 0.0055$ $0.20\pm 0.16$
+$3\times 10^{8}$ $0.065\pm 0.011$ $0.043\pm 0.0077$ $8.7\times 10^{-4}\pm 9.0\times 10^{-4}$ $0.030\pm 0.050$
+$3.25\times 10^{8}$ $0.076\pm 0.0093$ $0.048\pm 0.0010$ $0.0040\pm 0.0074$ $0.0019\pm 0.0034$
 Table 14: D-value for different $\sigma_{b}$.
-$\sigma_{b}$ d-value ($\dot{P}$) d-value ($P$)
-0.45 $0.0767\pm 0.00668$ $0.0513\pm 0.01258$
-0.48 $0.0621\pm 0.00880$ $0.0385\pm 0.01234$
-0.49 $0.0646\pm 0.00974$ $0.0304\pm 0.00501$
-0.5 $0.0557\pm 0.00740$ $0.0284\pm 0.00583$
-0.51 $0.056\pm 0.00735$ $0.0306\pm 0.00701$
-0.52 $0.0547\pm 0.00786$ $0.0358\pm 0.00579$
+$\sigma_{b}$ d-value ($\dot{P}$) d-value ($P$) p-value ($\dot{P}$) p-value ($P$)
+0.45 $0.077\pm 0.0067$ $0.051\pm 0.013$ $7.2\times 10^{-4}\pm 0.0011$ $0.0042\pm 0.0078$
+0.48 $0.062\pm 0.0088$ $0.038\pm 0.0123$ $0.0063\pm 0.0071$ $0.0063\pm 0.0071$
+0.49 $0.065\pm 0.0097$ $0.030\pm 0.0050$ $0.059\pm 0.057$ $0.16\pm 0.02437$
+0.5 $0.056\pm 0.0074$ $0.028\pm 0.0058$ $0.010\pm 0.014$ $0.22\pm 0.19$
+0.51 $0.056\pm 0.0073$ $0.031\pm 0.0070$ $0.0076\pm 0.0067$ $0.30\pm 0.18$
+Table 15: D-value for different $\sigma_{p}$.
+$\sigma_{p}$ d-value ($\dot{P}$) d-value ($P$) p-value ($\dot{P}$) p-value ($P$)
+0.4 $0.055\pm 0.0024$ $0.024\pm 0.0056$ $0.0036\pm 0.0050$ $0.57\pm 0.24$
+0.42 $0.054\pm 0.0055$ $0.028\pm 0.0051$ $0.0050\pm 0.0048$ $0.38\pm 0.23$
+0.45 $0.056\pm 0.0074$ $0.028\pm 0.0058$ $0.010\pm 0.014$ $0.22\pm 0.19$
+0.47 $0.053\pm 0.0049$ $0.034\pm 0.0092$ $0.0096\pm 0.015$ $0.26\pm 0.21$
+0.48 $0.053\pm 0.0045$ $0.037\pm 0.0084$ $0.0060\pm 0.0050$ $0.17\pm 0.18$
+0.52 $0.053\pm 0.0049$ $0.043\pm 0.011$ $0.0082\pm 0.012$ $0.11\pm 0.14$
 Table 16: D-value for different birth rate (BR).
-BR (in yr -1) d-value ($\dot{P}$) d-value ($P$)
-1/33 $0.0611\pm 0.01025$ $0.0313\pm 0.00636$
-1/37 $0.0565\pm 0.00851$ $0.0295\pm 0.00645$
-1/41 $0.0557\pm 0.00740$ $0.0284\pm 0.00583$
-1/45 $0.0585\pm 0.00510$ $0.0302\pm 0.00565$
-1/70 $0.0691\pm 0.00830$ $0.0342\pm 0.01165$
-1/150 $0.1165\pm 0.01564$ $0.0408\pm 0.00671$
+BR (in yr -1) d-value ($\dot{P}$) d-value ($P$) p-value ($\dot{P}$) p-value ($P$)
+1/33 $0.061\pm 0.010$ $0.031\pm 0.0064$ $0.013\pm 0.014$ $0.49\pm 0.17$
+1/37 $0.056\pm 0.0085$ $0.029\pm 0.0064$ $0.0086\pm 0.0095$ $0.48\pm 0.21$
+1/41 $0.056\pm 0.0074$ $0.028\pm 0.0058$ $0.010\pm 0.014$ $0.22\pm 0.19$
+1/45 $0.058\pm 0.0051$ $0.030\pm 0.0056$ $0.0055\pm 0.0042$ $0.18\pm 0.14$
+1/70 $0.069\pm 0.0083$ $0.034\pm 0.0116$ $4.3\times 10^{-4}\pm 0.0012$ $0.048\pm 0.074$
+1/150 $0.12\pm 0.016$ $0.041\pm 0.0067$ $5.9\times 10^{-12}\pm 1.2\times 10^{-11}$ $0.035\pm 0.091$

### v3

@@
 Extensive radio surveys, performed for instance by the Parkes and Arecibo radio-telescopes, so far 3700 pulsars have been discovered (Manchester et al.
-2005, see Table 1 listing the canonical pulsars).
+2005, see Table 1 listing the canonical pulsar population defined by isolated pulsars with periods longer than 20 msec and that are not magnetars).
 Although pulsars were first observed in radio, they were later found to be also bright in X-rays, optical and $\gamma$-rays.
@@
 The quantities $N_{\rm r}$, $N_{\rm g}$, and $N_{\rm rg}$ are the number of radio-only, gamma only, and radio-loud $\gamma$-ray pulsars, respectively.
-It should be noted that we excluded the binary pulsars, the pulsars in the Magellanic clouds and in globular clusters.
+It should be noted that we excluded the binary pulsars, the pulsars in the Magellanic clouds and in globular clusters, then we selected the pulsars with a magnetic field between $6\times 10^{6}$ T and $4.4\times 10^{9}$ T.
 The data have been taken from the ATNF catalogue from https://www.atnf.csiro.au/research/pulsar/psrcat/
@@
 (2022) only the canonical pulsar population in our Galaxy is reproduced.
-The canonical pulsar population is defined by isolated pulsars which are neither magnetars nor millisecond pulsars.
 While the study by Dirson et al.
@@
 $p(\log(B_{0}))=\frac{1}{\sigma_{b}\sqrt{2\pi}}e^{-(\log B_{0}-\log\bar{B})^{2}/(2\sigma_{b}^{2})}$ (3)
-$p(\log(P_{0}))=\frac{1}{\sigma_{p}\sqrt{2\pi}}e^{-(\log P_{0}-\log\bar{P})^{2}/(2\sigma_{p}^{2})}\ .$ (4)
+$p(\log(P_{0}))=\frac{1}{\sigma_{p}\sqrt{2\pi}}e^{-(\log P_{0}-\log\bar{P})^{2}/(2\sigma_{p}^{2})}$ (4)
 The distribution for the magnetic field is identical to the one used in other pulsar population synthesis like for instance Faucher-Giguère & Kaspi (2006), Gullon et al.
@@
 Thus, $\tau_{1}$ is another magnetic field decay time scale, corresponding to a certain magnetic field value $B_{1}$.
-Three different $\tau_{1}$ are considered in this work and are randomly chosen for each pulsar generated : $1.5\times 10^{5}$ yr or $3.5\times 10^{5}$ yr or $2.5\times 10^{6}$ yr with each of these timescales corresponds to a magnetic field value $B_{1}$ of $1\times 10^{8}$ T, $3\times 10^{8}$ T and $2\times 10^{9}$ T (these values of $\tau_{1}$ and $B_{1}$ are taken from Fig.
-10 of the work of Viganò et al.
-(2013)).
-Then the probabilities to get these values of $\tau_{1}$ and $B_{1}$ to be adopted for one pulsar are 0.23, 0.46 and 0.31 respectively (these probabilities were chosen as best educated guesses after several run with different values).
+Three different $\tau_{1}$ are considered in this work and are randomly chosen for each pulsar generated : $1.5\times 10^{5}$ yr or $3.5\times 10^{5}$ yr or $2.5\times 10^{6}$ yr with each of these timescales that corresponds to a magnetic field value $B_{1}$ of $1\times 10^{8}$ T, $3\times 10^{8}$ T and $2\times 10^{9}$ T.
+Then the probabilities to get these values of $\tau_{1}$ and $B_{1}$ to be adopted for one pulsar are 0.23, 0.46 and 0.31 respectively.
+These three probabilities actually approximate the probability density function for $\tau_{1}$ and $B_{1}$, and were already sufficient to reproduce the canonical population.
+Finding the complete distribution for $\tau_{1}$ and $B_{1}$ would require further investigations not needed in our study.
+Moreover, the distributions of $\tau_{1}$ and $B_{1}$ would physically represent the whole ensemble of trajectories possible in the $P-\dot{P}$ diagram for canonical pulsars as in Fig.
+10 of Viganò et al.
+(2013).
+Therefore, each pulsar has three possible evolutionary paths in the simulation, which is an empirical way to reproduce the $P-\dot{P}$ diagram in a better way than having only one evolutionary path.
+These particular choices for the discrete values were only guided by trial and errors.
 The value of $\alpha_{d}$ can be found in Table 5 of Sect.
@@
 Finally, we need to follow the particle motion within the Galactic potential. Each pulsar evolves in the gravitational potential $\Phi$ subject to an acceleration ${\ddot{x}}$ according to
-${\ddot{x}}=-\@vec{\nabla}\Phi\ .$ (13)
+${\ddot{x}}=-\@vec{\nabla}\Phi$ (13)
 Equation (13) is integrated numerically thanks to a Position Extended Forest Ruth-Like (PEFRL) algorithm (Young, 2014), a fourth order integration scheme, presented in Appendix A with a convergence and accuracy study.
@@
 $M_{b}$ $1.02\times 10^{10}\pm 6.3\times 10^{8}M_{\odot}$
-$M_{n}$ $4.10^{6}\pm 0.42\times 10^{6}M_{\odot}$
+$M_{n}$ $4\times 10^{6}\pm 0.42\times 10^{6}M_{\odot}$
 The potential for the bulge $\Phi_{b}$ and for the disk $\Phi_{d}$ were both chosen to have the form proposed by Miyamoto & Nagai (1975) which are typically used for models of the gravitational potential of the Milky Way.
@@
 2.4 The death line
-Several equations for the death line have been proposed in the literature (Chen & Ruderman, 1993; Zhang et al., 2000; Gil & Mitra, 2001; Mitra et al., 2019).
+Several equations for the death line have been proposed in the literature (Chen & Ruderman, 1993; Zhang et al., 2000; Gil & Mitra, 2001; Mitra et al., 2020).
 The most suitable one in our study is that of Mitra et al.
@@
 where $h_{\rm em}$ is the emission height, $P$ is the spin period of the pulsar and $c$ is the speed of light.
-The emission height is taken constant with an average value of $h_{\rm em}=3.10^{5}$ m, estimated from observations of a sample of pulsars by Weltevrede & Johnston (2008); Mitra (2017); Johnston & Karastergiou (2019); Johnston et al.
+The emission height is taken constant with an average value of $h_{\rm em}=3\times 10^{5}$ m, estimated from observations of a sample of pulsars by Weltevrede & Johnston (2008); Mitra (2017); Johnston & Karastergiou (2019); Johnston et al.
 (2023).
@@
 (2014) is used
-$\tau_{DM}=\frac{e^{2}\Delta f_{ch}\,DM}{\pi\,m_{e}\,c\,f^{3}}$ (30)
-$e$ is the electron charge, $m_{e}$ its mass, $\Delta f_{ch}$ the width of frequency of the instrument channel, $f$ the observing frequency and $DM$ the dispersion measure.
+$\tau_{DM}=\frac{e^{2}\Delta f_{ch}\,DM}{4\,\pi^{2}\,\epsilon_{0}\,m_{e}\,c\,f^{3}}=8.3\times 10^{15}\leavevmode\nobreak\ s\,\frac{\Delta f_{ch}}{f^{3}}\,DM$ (30)
+$e$ is the electron charge, $m_{e}$ its mass, $\Delta f_{ch}$ the width of frequency of the instrument channel in Hz, $f$ the observing frequency in Hz and $DM$ the dispersion measure in pc/cm3.
 Furthermore, in order to determine the influence of scattering by an inhomogeneous and turbulent ISM through $\tau_{scat}$, the empirical fit relationship from Krishnakumar et al.
@@
 $\tau_{scat}=3.6\times 10^{-9}DM^{2.2}(1+1.94\times 10^{-3}DM^{2})\ .$ (31)
-Both $\tau_{DM}$ and $\tau_{scat}$ depend on the dispersion measure, which is found for each pulsar by running the code of Yao et al.
+Both $\tau_{DM}$ and $\tau_{scat}$ are given in units of second (s) and depend on the dispersion measure DM (in units of pc/cm3), which is found for each pulsar by running the code of Yao et al.
 (2017) converting the distance of a pulsar into the dispersion measure thanks to their state-of-the art model of the Galactic electron density distribution.
@@
 In green, the period of the simulated pulsars with characteristic age greater than $10^{8}$ yr.
+Figure 3: Distribution of the observed age taken from the ATNF catalogue, along with the simulations without the implementation of the death line.
 Fig.
@@
 because we usually do not know the real ages of the observed pulsars.
-Figure 3: Distribution of the observed age taken from the ATNF catalogue, along with the simulations without the implementation of the death line.
+Figure 4: $P-\dot{P}$ diagram of the simulated population, along with the observations, without the implementation of the death line.
+Figure 5: Distribution of the observed period taken from the ATNF catalogue, along with the simulations with the death line implementation.
 In order to verify whether there is a correlation between old pulsars and those with a high spin period or not, Fig.
@@
 We conclude that canonical pulsars older than $10^{8}$ yr measured as their real age in our simulations should be discarded at the end of the simulation, since the characteristic age is overestimating the real age of pulsars.
-Indeed canonical pulsars older than this do not exist.
+Indeed, it appears that we are unable to detect canonical pulsars older than $10^{8}$ yr.
 The Galactic potential forbids many pulsars to escape from the Milky Way, they remain bound to the Galaxy, in contrast to the older pulsars found in Dirson et al.
@@
 4 but not implemented.
-Figure 4: $P-\dot{P}$ diagram of the simulated population, along with the observations, without the implementation of the death line.
 Given this excess of long period pulsars when neglecting the death line, we decided to implement it, in order to stick closer to reality.
 3.1.2 With the death line
-Figure 5: Distribution of the observed period taken from the ATNF catalogue, along with the simulations with the death line implementation.
 Adding a death line decreases the number of pulsars detected in radio compared to the situation without a death line, see Table 7.
@@
 3.2.2 Without a death line
+Figure 15: $P-\dot{P}$ diagram of the simulated population, along with the observations, without death line implementation, birth rate of 1/70 yr-1 for 1 million pulsars simulated.
+Figure 16: $P-\dot{P}$ diagram of the simulated population, along with the observations, without death line implementation, birth rate of 1/150 yr-1 for 1 million pulsars simulated.
 Simulation with 1 million simulated pulsars does not allow to generate pulsars older than $4.1\times 10^{7}$ yr.
@@
 Consequently, our attention turned towards exploring the impact of lower birth rates to ascertain if they provide a more congruous fit for both spin period and spin period derivative distributions.
-Figure 15: $P-\dot{P}$ diagram of the simulated population, along with the observations, without death line implementation, birth rate of 1/70 yr-1 for 1 million pulsars simulated.
-Figure 16: $P-\dot{P}$ diagram of the simulated population, along with the observations, without death line implementation, birth rate of 1/150 yr-1 for 1 million pulsars simulated.
 Running simulations with a birth rate of $1/70$ yr-1 or a birth rate of $1/150$ yr-1 (as a consequence, increasing the oldest simulated pulsar), conspicuously revealed an insufficient number of detected pulsars in both Fig.
@@
 However if we smoothly change $F_{\rm min}$ from $4\times 10^{-15}$ to $16\times 10^{-15}$ (the two $F_{\rm min}$ values used in this work), the bi-modality would almost disappear.
+Finally, we compare the simulated and observed $\gamma$-ray light-curve peak separation $\Delta$ in Fig. 22. The separation $\Delta$ is computed according to Pétri (2011) by
+$\cos\left(\pi\,\Delta\right)=|\cot\xi\cot\alpha|$ (37)
+where $\xi$ is the angle between the line of sight and the rotation axis and $\alpha$ the inclination angle.
+The distributions are normalized in order to obtain probability distribution functions (p.d.f) for both observations and simulations.
+The normalization is performed by dividing each distribution function by its total number of pulsars, observed or simulated.
+Moreover, $\Delta$ is constrained to the interval between 0 and 0.5, because we consider that peak separations $\Delta$ larger than 0.5 can be scaled to $\Delta=1-\Delta$ by a phase shift, since the signal is periodic and represents the phase separation between two peaks and the definition of the first or main peak is arbitrary.
+Fig.
+22 displays a remarkable similarity between the observed and simulated distributions of the $\gamma$-ray peak separation, meaning the striped wind model captures very well this geometric feature.
 The higher number of detected pulsars in the simulations could be related to the unidentified sources in the Fourth Fermi-LAT catalogue of $\gamma$-ray sources (4FGL).
@@
 Figure 21: Distribution of the $\gamma$-ray flux of the simulated population along with the observations from the 3PC catalogue.
+Figure 22: Distribution of the $\gamma$-ray peak separation of the simulated population along with the observations from the 3PC catalogue.
 To conclude the part about $\gamma$-ray pulsars, we predict the number of pulsars that would be detected by a future instrument, ten times more sensitive than Fermi/LAT.
@@
 The $\gamma$-ray detected pulsar positions in the X-Y plane of the Galaxy are shown in Fig.
-22.
+23.
 In this simulation, 2492 $\gamma$-ray pulsars are detected, in green, being $\gamma$-ray only or radio-loud $\gamma$-ray pulsars, among them only 350 (really close to the number of $\gamma$-ray and radio/$\gamma$-ray pulsars detected in Subsection 3.2) would have been detected if the same sensitivity as before was kept, in red in Fig.
-22.
+23.
 Therefore we detect 7 times more $\gamma$-ray pulsars compared to the number found with the previous sensitivity.
 Furthermore, the other 2142 pulsars, in green in Fig.
-22, detected with this increased sensitivity highlight several interesting features.
+23, detected with this increased sensitivity highlight several interesting features.
 We obviously detect more pulsars close to Earth, and also many more in the centre of the Galaxy.
@@
 Nonetheless, simulating 1 million pulsars allowed to show that with or without a death line the results were alike and therefore, in the case where the oldest simulated pulsar is $4.1\times 10^{7}$ yr (its real age), a death line is not necessary to explain the observations.
-In this paper we showed two approaches to reproduce the observed canonical pulsar population, however when we consider the approach where the oldest pulsar simulated is only $4.1\times 10^{7}$ yr we state that no older pulsars might exist in the Galaxy, while the Milky Way is very much older than this.
-There is apparently no reason not to simulate pulsars that are older than this, and if there is one we do not know it, they become harder to detect because of their wide pulse profile with their age, but some are still detectable.
+In this paper we showed two approaches to reproduce the observed canonical pulsar population, however when we consider the approach where the oldest pulsar simulated is only $4.1\times 10^{7}$ yr we state that no older pulsars might be detected with our current telescopes in the Galaxy, while the Milky Way is very much older than this and therefore has been forming pulsars since its beginning.
+There is apparently no reason not to simulate pulsars that are older than $10^{8}$ yr, and if there is one we do not know it, they become harder to detect because of their wide pulse profile with their age, but some could still be detectable.
 If this situation is the reality then it means that the death line is not necessary in pulsar population synthesis to explain observations.
@@
 The trajectory obtained after the integration with the PEFRL algorithm is shown in Fig.
-23.
+24.
 This trajectory shows that the pulsar, with these initial conditions, remains bound to the Galaxy.
@@
 The goal is here is to show the results with a constant magnetic field in order to see which model is better between the decaying magnetic field and the constant magnetic field.
-Hence, this simulation was run with 1 million pulsar simulated, without a death line and without a decaying of the magnetic field.
-The KS test gave p-values of $1.1\times 10^{-14}$ and $1.5\times 10^{-76}$, therefore the null hypothesis is rejected for both $P$ and $\dot{P}$.
+Hence, this simulation was run with 30 million pulsar simulated (meaning the oldest pulsar simulated is $1.23\times 10^{9}$ yr), with a death line and without a decaying magnetic field.
+The KS test gave p-values of $1.03\times 10^{-22}$ and $3.38\times 10^{-51}$ for $P$ and $\dot{P}$ respectively, therefore the null hypothesis is rejected for both $P$ and $\dot{P}$.
+Here we simulate more pulsars than in the main part of the paper, since if we simulate with the same number of pulsars, it seems the whole $P-\dot{P}$ diagram is not covered.
+Therefore if a constant magnetic field was the best way to reproduce the canonical population, it means we would detect much older pulsars than with the magnetic field decay prescription.
 The parameters used are in Table 5.
@@
 As can be seen in Fig.
-25, too many pulsars from the simulation appear on the top right of the diagram and almost none at the bottom.
+26, too many pulsars from the simulation appear on the top right of the diagram and almost none at the bottom.
 In Table 11, we note a total of pulsars which is smaller from what we found for all of our previous results in Sect.
@@
 log($\dot{E}$) (W) $N_{tot}$ $N_{r}$ $N_{g}$ $N_{rg}$
-¿31 14 0 10 4
-¿28 157 6 77 74
-Total 1328 974 176 178
+¿31 42 0 29 13
+¿28 213 10 127 76
+Total 1473 1075 215 183
 Appendix C Influence of the different parameters